### v3

# Peter Reinhard Hansen§    and   Chen Tong

Abstract
We introduce a novel method for obtaining a wide variety of moments of a random variable with a well-defined moment-generating function (MGF). We derive new expressions for fractional moments and fractional absolute moments, both central and non-central moments. The new moment expressions are relatively simple integrals that involve the MGF, but do not require its derivatives. We label the new method CMGF because it uses a complex extension of the MGF and can be used to obtain complex moments. We illustrate the new method with three applications where the MGF is available in closed-form, while the corresponding densities and the derivatives of the MGF are either unavailable or very difficult to obtain.
Keywords: Moments, Fractional Moments, Moment-Generating Function.
JEL Classification: C02, C40, C65
1 Introduction
Moments of random variables, including conditional moments, are central to much of economics, finance, and econometrics. In this paper, we introduce a new method for obtaining moments of random variables with a well-defined moment-generating function (MGF). The proposed method is general, computationally efficient, and can offer solutions to problems where standard methods are inadequate.
It is well known that the $k$-th integer moment of a random variable $X$, with MGF, $M_{X}(s)=\mathbb{E}[e^{sX}]$, is given by $\mathbb{E}[X^{k}]=M_{X}^{(k)}(0)$, where $M_{X}^{(k)}(s)$ is the $k$-th derivative of $M_{X}(s)$. Non-integer moments and fractional absolute moments, $\mathbb{E}|X^{r}|$, $r\in\mathbb{R},$ have more complicated expressions, typically integrals that involve derivatives of the MGF or derivatives of the characteristic function (CF). [^1] Kawata, (1972) provides an expression for $\mathbb{E}|X|^{r}$ that involves several derivatives of the CF and Laue, (1980) provides an expression based on fractional derivatives, see also Wolfe, (1975), Samko et al., (1993), Matsui and Pawlas, (2016), and Tomovski et al., (2022). For fractional moments of non-negative variables, similar expressions were derived in Cressie et al., (1981), Cressie and Borkent, (1986), Jones, (1987), and Schürger, (2002). In this paper, we derive novel integral expressions for computing fractional moments, including fractional absolute moments, and fractional central moments. Unlike existing expressions, the new expressions do not involve derivatives (of the MGF or the CF). Thus, the new expressions are particularly useful when derivatives are prohibitively difficult to obtain. This is often the case in dynamic models, where the conditional MGF is defined from recursive non-linear expressions. The new method is labelled CMGF because it relies on a complex extension of the MGF, and provides expressions for complex moments, $\mathbb{E}|X|^{r}$, $r\in\mathbb{C}$. Although complex moments are not commonly used in econometrics, this generalization can be included with minimal adaptation.  [^2] Complex moments are commonly used other fields, including statistical mechanics, number theory, and quantum physics. In part, because complex moments are related to Mellin transforms. The new expressions can also be used for establishing new identities, such as the expression for the reciprocal Gamma function, which we infer from the absolute moment of a Gaussian random variable.
The new CMGF expressions involve integrals, and while these cannot be evaluated analytically in most cases, they facilitate new ways to compute moments numerically, which are fast and accurate. We demonstrate this with the normal-inverse Gaussian distribution. We also show that the CMGF method offers fast and accurate moments in dynamic models, where we are not aware of good alternative methods. Specifically, we use the CMGF method to compute moments of cumulative returns in the Heston-Nandi GARCH (HNG) model, to compute moments of realized volatilities in the Heterogeneous Autoregressive Gamma (HARG) model, and to compute conditional moments in the Autoregressive Poisson Model.
The remainder of this paper is organized as follows. We present the new moment expressions in Section 2. Theorem 1 has the expression for absolute fractional moments for general random variables, Theorem 2 has the expression for fractional moments of non-negative random variables, and Theorem 3 presents a new expression for integer moments. As an example, we apply CMGF to the normal-inverse Gaussian distribution. Section 3 presents three applications of the new expressions in dynamic models, and Section 4 concludes.
2 A New Method for Computing Moments (CMGF)
We will use the following notation: $\mathbb{N}$ denotes the positive integers and $\mathbb{N}_{0}=\{0\}\cup\mathbb{N}$. We use $\mathcal{N}$ to denote a neighborhood about zero, for which $(-\delta,\delta)\subset\mathcal{N}$ for some $\delta>0$, and the positive part of $\mathcal{N}$ is denoted $\mathcal{N}_{+}\equiv\{s\in\mathcal{N}:s>0\}$. We use $z=s+it$ and $\zeta=\sigma+i\tau$ to represent complex variables, where $s=\operatorname{Re}(z)$ and $t=\operatorname{Im}(z)$, and similarly for $\zeta$.
To avoid possible confusion, we use $\mathbb{E}[X^{k}]$ to denotes integer moments, $k\in\mathbb{N}$, whereas general moments (including fractional and complex moments) are represented with $\mathbb{E}[X^{r}]$, $r\in\mathbb{C}$.
Our results are applicable to any random variable with a well-defined MGF, as characterized in the following assumption.
Assumption 1.
$X$ is a real random variable with
$M_{X}(s)=\mathbb{E}[e^{sX}]<\infty,\qquad\text{for all}\quad s\in\mathcal{N},$
where $\mathcal{N}$ is a neighborhood about zero.
The MGF is typically used as a real function with a real argument, such that both $s$ and $M_{X}(s)$ take values in $\mathbb{R}.$ In this paper, we establish our results by means of complex arguments, $z\in\mathbb{C},$ such that $M_{X}(z)$ also takes values in $\mathbb{C}$.
To gain some intuition for the complex extension of the MGF, we observe that
$\psi(s,t)\equiv M_{X}(s+it)=\mathbb{E}[e^{(s+it)X}]\in\mathbb{C},$
nests both the standard moment-generating function, $\psi(s,0)=M_{X}(s)$, and the characteristic function, $\psi(0,t)=\varphi_{X}(t)\equiv\mathbb{E}[e^{itX}]$, as special cases. The complex valued MGF, $M_{X}(z)$, is the bilateral Laplace–Stieltjes transform (apart from the sign of $z$), and for continuous random variables it is similar to the Laplace transform.  [^3] The Laplace transform of the density function, $f_{X}(x)$, is $\int_{0}^{\infty}e^{-zx}f_{X}(x)\mathrm{d}x$, and it differs from $M_{X}(z)$ in terms of the domain of integration and the sign of the exponent. Moreover, $M_{X}(z)$ is obviously also applicable to discrete distributions and distributions that are mixtures of continuous and discrete distributions.
The results are based on the following identities.
Lemma 1.
Let $x\in\mathbb{R}$ and $z=s+it\in\mathbb{C}$, where $s>0$ is an arbitrary positive constant.
$|x|^{r}=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}+e^{-zx}}{z^{r+1}}\mathrm{d}t,\qquad\operatorname{Re}(r)>-1,$
for any $x\neq 0$, and this identity also holds for $x=0$ if $\operatorname{Re}(r)>0$. For $x>0$ we have
$x^{r}=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}t,\qquad\operatorname{Re}(r)>-1.$
Similarly,
$x^{k}$ $=$ $\frac{k!}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}+\left(-1\right)^{k}e^{-zx}}{z^{k+1}}\mathrm{d}t,\qquad k\in\mathbb{N}_{0},$
for any $x\neq 0$, and this identity also holds for $x=0$ if $k\in\mathbb{N}$.
The case, $x=0$, arises as a special case in Lemma 1, requiring results to include a qualification for distributions with positive mass at zero.
Our first result is the following expression for fractional absolute (possibly complex) moments.
Theorem 1 (Absolute moments).
Suppose $X$ satisfies Assumption 1, $\xi\in\mathbb{R}$, and $z=s+it$ with $s\in\mathcal{N}_{+}$. Then
$\mathbb{E}|X-\xi|^{r}$ $=$ $\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-\xi z}M_{X}(z)+e^{\xi z}M_{X}(-z)}{z^{r+1}}\mathrm{d}t,$ (1)
for $\operatorname{Re}(r)\geq 0$, and if $\Pr(X=\xi)=0$, then the identity holds for all $\operatorname{Re}(r)>-1$.
If $r$ is real, then (1) is equivalent to
$\mathbb{E}|X-\xi|^{r}=\frac{\Gamma(r+1)}{\pi}\int_{0}^{\infty}\operatorname{Re}\left[\frac{e^{-\xi z}M_{X}(z)+e^{\xi z}M_{X}(-z)}{z^{r+1}}\right]\mathrm{d}t.$ (2)
The result obviously covers regular and central absolute moments by setting $\xi=0$ for regular moments and $\xi=\mathbb{E}X$ for central moments. Note that the expression remains invariant to the specific value of $s$, so long as $s\in\mathcal{N}_{+}$. In our experience, the choice of $s$ has no practical impact on moment evaluation. Although analytical integration may often be impractical, (1) offers a novel method to evaluate moments numerically. For real moments, $r\in\mathbb{R}$, it is our experience that it is computationally advantageous to use (2) rather than (1).
Interestingly, Theorem 1 can be used to obtain new identities, by equating (1) with an existing expression for absolute moments. For instance, the absolute moment for standard normal random variable, $X\sim N(0,1)$, is $\mathbb{E}|X|^{r}=\frac{\Gamma(\frac{r+1}{2})}{\sqrt{\pi}}2^{r/2}$, for $r>-1$, and equating this with (1) yields the following expression for the reciprocal gamma function
$\frac{1}{\Gamma(\frac{r}{2}+1)}=\frac{2^{r/2}}{\pi}\int_{-\infty}^{+\infty}\frac{e^{z^{2}/2}}{z^{r+1}}\mathrm{d}t,\qquad r>-1,$ (3)
for $z=s+it$, where $s>0$ is an arbitrary positive constant. While we can only infer that (3) holds for real $r>-1$, the identity actually holds for any $r\in\mathbb{C}$, see Hansen and Tong, (2024).
For non-negative random variables we have the following simpler expression for fractional moments.
Theorem 2 (Moments of non-negative random variables).
Let $X$ satisfies Assumption 1 and $\Pr(X\geq\xi)=1$ for some $\xi\in\mathbb{R}$. For any $r\in\mathbb{C}$ with $\operatorname{Re}(r)\geq 0$, we have,
$\mathbb{E}[(X-\xi)^{r}]=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-\xi z}M_{X}(z)}{z^{r+1}}\mathrm{d}t,$ (4)
where $z=s+it$ with $s\in\mathcal{N}_{+}$. If $X$ is strictly greater than $\xi$, i.e. $\Pr(X>\xi)=1$, then the identity holds for $\operatorname{Re}(r)>-1$.
If $r$ is real, then (4) is equivalent to
$\mathbb{E}[(X-\xi)^{r}]=\frac{\Gamma(r+1)}{\pi}\int_{0}^{\infty}{\rm Re}\left[\frac{e^{-\xi z}M_{X}(z)}{z^{r+1}}\right]\mathrm{d}t.$ (5)
As an illustration, we apply (4) to an exponentially distributed random variable. This example serves as an illustration because integer moments, $k\in\mathbb{N}$, are easy to obtain from the derivatives, $M_{X}^{(k)}(s)$, and other moments can be obtained by evaluating $\int_{0}^{\infty}x^{r}\lambda e^{\lambda x}\mathrm{d}x$ directly.
Example 1.
For an exponentially distributed random variable, with parameter $\lambda>0$, $X\sim\operatorname{Exp}(\lambda)$, we have $M_{X}(z)=\frac{\lambda}{\lambda-z}$, with $\mathcal{N}=(-\infty,\lambda)$. So, the integral in (5) is $\int_{0}^{\infty}{\rm Re}\left[\frac{\lambda}{\lambda-z}\frac{1}{z^{r+1}}\right]\mathrm{d}t$. For an integer moment, $k\in\mathbb{N}$, this integral equals
$\int_{0}^{\infty}{\rm Re}\left[\lambda^{-k}\tfrac{1}{\lambda-z}+\lambda^{-k}\tfrac{1}{z}+\sum_{j=2}^{k+1}\lambda^{-k+j-1}\tfrac{1}{z^{j}}\right]\mathrm{d}t$ $=\lambda^{-k}\int_{0}^{\infty}{\rm Re}\left[\tfrac{1}{\lambda-z}+\tfrac{1}{z}\right]\mathrm{d}t$
$=\lambda^{-k}\int_{0}^{\infty}{\rm Re}\left[\tfrac{\lambda-z^{\ast}}{(\lambda-z^{\ast})(\lambda-z)}+\tfrac{z^{\ast}}{z^{\ast}z}\right]\mathrm{d}t,$
because $\int_{0}^{\infty}{\rm Re}\left[z^{-j}\right]\mathrm{d}t=0$ for $j>1$. Substituting $z=s+it$ in the last expression gives us
$\lambda^{-k}\int_{0}^{\infty}\operatorname{Re}\left[\tfrac{\lambda-s+it}{(\lambda-s)^{2}-(it)^{2}}+\tfrac{s-it}{s^{2}-(it)^{2}}\right]\mathrm{d}t$ $=$ $\lambda^{-k}\int_{0}^{\infty}\left[\tfrac{\lambda-s}{(\lambda-s)^{2}+t^{2}}+\tfrac{s}{s^{2}+t^{2}}\right]\mathrm{d}t$
$=$ $\lambda^{-k}\left[\left.\arctan(\tfrac{t}{\lambda-s})+\arctan(\tfrac{t}{s})\right|_{0}^{\infty}\right]=\pi\lambda^{-k}.$
Thus that $\mathbb{E}[X^{k}]=\lambda^{-k}k!$ as expected. Deriving non-integer moments, $r\notin\mathbb{N}$, can be derived using contour integrals and the Cauchy Integral Theorem. These are the techniques used extensively to prove the main results in this paper.
For random variables whose support may include negative numbers, we have the following result for integer moments.
Theorem 3.
Let $X$ satisfy Assumption 1 and let $\xi\in\mathbb{R}$. For $k\in\mathbb{N}$ we have
$\mathbb{E}[(X-\xi)^{k}]$ $=$ $\frac{k!}{\pi}\int_{0}^{\infty}\operatorname{Re}\left[\frac{e^{-\xi z}M_{X}(z)+(-1)^{k}e^{\xi z}M_{X}(-z)}{z^{k+1}}\right]\mathrm{d}t,$ (6)
where $z=s+it$ with $s\in\mathcal{N}_{+}$.
Because $k$ is real, we have expressed the integral on the form, $\frac{k!}{\pi}\int_{0}^{\infty}\operatorname{Re}[\cdot]\mathrm{d}t$, which tends to be simpler to evaluate numerically than the equivalent expression using the form, $\frac{k!}{2\pi}\int_{-\infty}^{+\infty}[\cdot]\mathrm{d}t$.
Remark.
In some situations, Theorems 1, 2, and 3, provide different expressions for the same moments. For instance, (4) and (1) must agree for a non-negative random variable, because $\mathbb{E}|X|^{r}=\mathbb{E}[X^{r}]$ in this case, and for integer moments, these expressions must also agree with (6). This is indeed the case, because the additional term in (1) and the additional term in (6) are both zero for non-negative variables. This is an implication of Lemma A.2, which establishes that $\int_{-\infty}^{\infty}\frac{e^{-\left(s+it\right)x}}{\left(s+it\right)^{r+1}}\mathrm{d}t=0$ for all $x>0$ and all $s>0$.
We can combine Theorems 1 and 3 to compute odd moments over tail events.
Corollary 1 (Tail Moments).
Let $X$ satisfy Assumption 1, $\xi\in\mathbb{R}$, and let $k$ be odd number. Then
$\mathbb{E}[(X-\xi)^{k}1_{\{X>\xi\}}]$ $=$ $+\frac{k!}{\pi}\int_{0}^{\infty}{\rm Re}\left[\frac{e^{-\xi z}M_{X}(z)}{z^{k+1}}\right]\mathrm{d}t,$
$\mathbb{E}[(X-\xi)^{k}1_{\{X<\xi\}}]$ $=$ $-\frac{k!}{\pi}\int_{0}^{\infty}{\rm Re}\left[\frac{e^{-\xi z}M_{X}(z)}{z^{k+1}}\right]\mathrm{d}t,$
where $z=s+it$ for any $s\in\mathcal{N}_{+}$.
We can also compute the cdf, quantiles, and expected shortfall using the new moment expressions. We can obtain the cdf for a continuous random variable, $X$, from the MGF by Gil-Pelaez inversion,
$F_{X}(x)=\frac{1}{2}-\frac{1}{\pi}\int_{0}^{\infty}{\rm Re}\left[\frac{M_{X}(it)e^{-itx}}{it}\right]\mathrm{d}t,$
where we used that $M_{X}(it)$ is the characteristic function for $X$. Now the $\alpha$-quantile is given by $\xi_{\alpha}\equiv F_{X}^{-1}(\alpha)$, and we can use Corollary 1 (with $k=1$) to compute,
$\operatorname{ES}_{\alpha}(X)=-\frac{1}{\alpha}\mathbb{E}[X1_{\{X<\xi_{\alpha}\}}]=-\frac{1}{\alpha}\mathbb{E}[(X-\xi_{\alpha})1_{\{X<\xi_{\alpha}\}}]-\xi_{\alpha},$
which is known as the expected shortfall.
2.1 Example: Moments of NIG Distribution
To illustrate the new method, we use Theorem 1 to compute fractional moments of the normal-inverse Gaussian (NIG) distribution. The CMGF method is significantly faster and more accurate than both simulation-based methods and direct integration using the NIG density.
The NIG distribution was introduced in Barndorff-Nielsen, (1978) and has four parameters, $\lambda$ (location), $\delta$ (scale), $\alpha$ (tail heaviness), and $\beta$ (asymmetry). It has density
$f(x)=\tfrac{\alpha\delta}{\pi\sqrt{\delta^{2}+(x-\lambda)^{2}}}K_{1}\left(\alpha\sqrt{\delta^{2}+(x-\lambda)^{2}}\right)e^{\delta\gamma+\beta(x-\lambda)},\qquad x\in\mathbb{R}$
where $\gamma=\sqrt{\alpha^{2}-\beta^{2}}$ and $K_{1}(\cdot)$ is the modified Bessel function of the second kind, and the MGF is given by $M_{X}(z)=\exp\left(\lambda z+\delta\left[\gamma-\sqrt{\alpha^{2}-(\beta-z)^{2}}\right]\right),$ for all $z\in\mathbb{C}$. We consider two standardized NIG distribution (zero mean and unit variance), which can be characterized by the two parameters, $(\xi,\chi)$, for $10\leq|\chi|<\xi<$ 1, where $\lambda=-\tfrac{\chi}{\xi}\zeta$, $\delta=\xi^{2}\sqrt{\xi^{2}-\chi^{2}}\zeta$, $\alpha=\xi\zeta$, and $\beta=\chi\zeta$, with $\zeta=\sqrt{1-\xi^{2}}/(\xi^{2}-\chi^{2})$, see Barndorff-Nielsen et al., (1985). Specifically we consider $(\xi,\chi)=(1/2,-1/3)$ and $(\xi,\chi)=(1/8,-1/16)$, the corresponding densities are shown in the left panel of Figure 1. The corresponding MGFs, $t\mapsto\operatorname{MGF}(s+it)$ are shown in the right panel for $s=1$ – solid lines for the real part and dashed lines for the complex part.
[…]
We compute the absolute moments for $-0.85\leq r\leq 4.2$ for both distributions using Theorem 1 and the moments are shown in the upper panel of Figure 2. The solid black dot is the second (absolute) moment, which is known to be one for these standardized distributions, and the solid blue and red dots represent the fourth (absolute) moment, which are known to equal $3(1+4\chi^{2})/(1-\xi^{2})$. The values provided by the CMGF method are very accurate. The first 12 digits of the fourth moments are correct using a simple implementation in Julia.
Figure 2: Absolute moments of the standardized NIG distribution with $(\xi,\chi)=(1/2,-1/3)$ (blue lines) and $(\xi,\chi)=(1/8,-1/16)$ (red lines). These are shown in the left panels, where dots are represent the known moments for $r=2$ and $r=4$. The middle panel is a snippet of the left panel which can be compared to simulation-based moments for 15 values of $r$. For each $r$ we computed $\mathbb{E}|X|^{r}$ 100 times as represented by the x-crosses. Each of these simulation-based moments are based on $N=$1,000,000 draws of NIG distribution. The right panel shows how many decimal places are accurate with a one standard deviation simulation error.
Additionally, we compute the absolute moments using $N=1,000,000$ independent draws from the two NIG distributions. The simulation based moment is given by $\hat{\mu}_{N}^{r}=\frac{1}{N}\sum_{j=1}^{N}|X_{i}|^{r}$ and we estimate the $\sigma^{2}(r)=\operatorname{var}(|X_{j}|^{r})$ with 100 million draws. The accuracy of a simulated moment can be computed from its standard error. Assuming an error equal to one standard deviation, then the number of accurate decimal places is given by $-\log_{10}(\sigma(r)/\sqrt{N})$. [^4] For instance, an error where the first three decimal places are zero, say $0.00099$, will map to $-\log_{10}(0.99\cdot 10^{-3})\simeq 3.004$. We have plotted this number for $N=10^{6}$ in the lower left panel of Figure 2. The simulation-based moments based on $N=1,000,000$ draws from the distribution are far less accurate. At best, the first few decimal places are accurate, and the accuracy deteriorates rapidly as $r$ increases. In the lower right panels we have shown the moments based on the CMGF method for $3.8\leq r\leq 4.2$ represented by the solid blue line. The $\times$-crosses are simulation-based moment estimates, each based on $N=1,000,000$ random draws.
Table 1: CMGF Computation Time of $\mathbb{E}|X|^{r}$ for $X\sim\operatorname{NIG}$
$r$ CMGF Integrate with density Simulations $(N=10^{6})$
-0.5 28.3 712.9 397,666
0.5 29.2 244.0 395,837
1.0 20.3 199.1 299,232
1.5 29.3 175.0 399,173
2.0 20.4 108.3 298,635
2.5 24.5 149.3 399,182
3.0 17.0 122.5 291,924
3.5 24.5 114.7 402,552
4.0 17.6 100.3 318,081
Note: Computation time (in microseconds) for evaluating $\mathbb{E}|X|^{r}$ using three methods: the CMGF method of Theorem 1, by numerical integration of $\int_{-\infty}^{\infty}|x|^{r}f(x)\mathrm{d}x$, and by simulations, where we generate $X_{1},\ldots,X_{N}$ independent and identically NIG distributed, with $N=$1,000,000 and take the average of $|X_{i}|^{r}$. Both CMGF and the density-based method use numerical integration with the same tolerance threshold. Computation times are evaluated with BenchmarkTools.jl for Julia, see Chen and Revels, (2016). Computation were done with Julia v1.11.0, see Bezanson et al., (2017), on a MacBook Pro M1 Max with 32 GM memory.
The CMGF method is not only far more accurate in this example, it is also much faster than the simulation-based approach. This is shown in Table 1 where we report the computation time obtained with a Julia implementation. In this case the density is readily available, and we also compute the moment by evaluating $\int_{-\infty}^{\infty}|x|^{r}f(x)\mathrm{d}x$ using the same numerical integration method used to evaluate the integral in Theorem 1. The CMGF method is between 5 and 25 times faster than the standard approach, and the CMGF method is more than 10,000 times faster than averaging (and generating) $N=10^{6}$ independent draws from the NIG distribution. We have also compared the methods using a standard normal distribution, for which the true moment is known to be, $\mathbb{E}|X|^{r}=\frac{\Gamma(\frac{r+1}{2})}{\sqrt{\pi}}2^{r/2}$, for all $r>-1$. For the Gaussian case the CMGF is also thousands of times faster than simulation methods using 1 million draws and far more accurate. For $|r|<1$, the CMGF is also faster and more accurate than numerical integration based on the density (even if we take advantage of the density being symmetric about zero). For moments larger than one, integration with the density is about twice as fast as CMGF, after taking advantage of the symmetric density.
A multivariate extension is beyond the scope of this paper. However, some simple cross-moments, such as $\mathbb{E}[X_{1}X_{2}]$ and $\mathbb{E}[X_{1}X_{2}^{2}]$, are easy to obtain from univariate results. Let $(X_{1},X_{2})$ have MGF, $M_{X}(t_{1},t_{2})$, such that $M_{X_{1}+X_{2}}(t)=M_{X}(t,t)$, $M_{X_{1}-X_{2}}(t)=M_{X}(t,-t)$, $M_{X_{1}}(t)=M_{X}(t,0)$, and $M_{X_{2}}(t)=M_{X}(0,t)$, from which we obtain $\mathbb{E}[(X_{1}+X_{2})^{k}]$, $\mathbb{E}[(X_{1}-X_{2})^{k}]$, etc. by Theorem 3, and the two cross moments are given by $\mathbb{E}[X_{1}X_{2}]=\frac{1}{2}\left(\mathbb{E}[(X_{1}+X_{2})^{2}]-\mathbb{E}[X_{1}^{2}]\mathbb{E}[X_{2}^{2}]\right)$ and $\mathbb{E}[X_{1}X_{2}^{2}]=\frac{1}{6}\left(\mathbb{E}[(X_{1}-X_{2})^{3}]+\mathbb{E}[(X_{1}+X_{2})^{3}]-2\mathbb{E}[X_{1}^{3}]\right)$, respectively.
In the next Section, we compute conditional moments in dynamic models with the CMGF methods. The appropriate density is unknown in most of these problems, and this makes the CMGF method particularly useful.
3 Applications to Dynamic Models
We consider three applications that demonstrate the usefulness of the new CMGF expressions. In the first application, we use Theorems 1 and 3 to obtain moments for cumulative returns from a Heston-Nandi GARCH model. The second application uses Theorem 2 to compute fractional moments of realized volatility that follow an Autoregressive Gamma process. The third application also applies Theorem 2 to obtain fractional moments, but in this case for a discrete distribution that has probability mass at zero.
A common feature of these three applications is that the relevant MGF is available in closed form, while the corresponding density and its derivatives are extremely difficult, if not impossible, to obtain. Although simulation-based methods are always an option, they will tend to be much slower than numerical integration of the new analytical expressions. We will also implement simulation-based methods to compare such with the new moment expressions, and compare their relative computational burden.
We use the following notation: $\mathbb{E}_{t}(\cdot)=\mathbb{E}(\cdot|\mathcal{F}_{t})$ is the conditional expectation where $\{\mathcal{F}_{t}\}$ is the natural filtration. We will typically condition on $\mathcal{F}_{T}$, and use $M_{X_{T+H}|T}(z)\equiv\mathbb{E}[\exp(X_{T+H}z)|\mathcal{F}_{T}]$ to denote the conditional MGF for the random variable $X_{T+H}\in\mathcal{F}_{T+H}$.
3.1 Moments of Cumulative Returns in a GARCH Model
The Heston-Nandi GARCH (HNG) model by Heston and Nandi, (2000) is given by
$r_{t+1}$ $=r_{f}+\left(\lambda-\tfrac{1}{2}\right)h_{t+1}+\sqrt{h_{t+1}}z_{t+1},\qquad\text{with}\quad z_{t}\sim iidN(0,1),$ (7)
$h_{t+1}$ $=\omega+\beta h_{t}+\alpha\left(z_{t}-\gamma\sqrt{h_{t}}\right)^{2},$
where $r_{f}$ is the risk free rate; $\lambda$ is the equity risk premium; and $h_{t+1}={\rm var}\left(R_{t+1}|\mathcal{F}_{t}\right)$ is the conditional variance of daily log-returns $r_{t+1}=\log\left(S_{t+1}/S_{t}\right)$.
Among the many variations of GARCH models, the HNG model stands out as having an analytical MGF for cumulative returns, $R_{T,H}=\sum_{t=T+1}^{T+H}r_{t}$. The underlying reason is that its dynamic structure is carefully crafted for the purpose of yielding closed-form option pricing formulae, and option pricing formulae depend on the properties of cumulative returns, $R_{T,H}=\sum_{t=T+1}^{T+H}r_{t}$. The conditional density function for $R_{T,H}$ does not have a known analytical form, but the corresponding MGF is available in closed form [^5] We follow standard terminology and label an expression as “closed form” if it can be expressed in terms of elementary mathematical functions and operations, which is the case for our recursive expressions. from simple recursive expressions, as stated below.
Proposition 1.
Let $r_{t}$, $t=1,2,\ldots$, be given by (7) and define cumulative returns, $R_{T,H}=r_{T+1}+\cdots+r_{t+H}$. Then the conditional MFG has the affine form,
$M_{R_{T,H}|T}(z)=\exp\left(A(H,z)+B(H,z)h_{T+1}\right),$ (8)
which is well-defined for $z\in\{\zeta\in\mathbb{C}:$$B(h,\operatorname{Re}(\zeta))<\frac{1}{2\alpha}$, for all $h=1,\ldots,H\}$, where $A(H,z)$ and $B(H,z)$ are given from the recursions,
$A(h+1,z)$ $=A(h,z)+zr_{f}+B(h,z)\omega-\frac{1}{2}\log\left(1-2B(h,z)\alpha\right),$
$B(h+1,z)$ $=z(\lambda-\tfrac{1}{2})+B(h,z)\left(\beta+\alpha\gamma^{2}\right)+\frac{\left(z-2\alpha\gamma B(h,z)\right)^{2}}{2\left(1-2\alpha B(h,z)\right)},$
with initial values, $A(1,z)=zr_{f}$ and $B\left(1,z\right)=z(\lambda-\frac{1}{2})+\frac{z^{2}}{2}$.
Obtaining the moments of cumulative return by way of the derivatives of the MGF is nearly impossible, especially for higher moments and for cumulative returns over many periods (large $H$). Instead, we can compute the moments by Theorems 1 and 3, where the former also enables us to compute fractional absolute moments.
We will illustrate this with a simulation design that is based on real data. We estimate the HNG model using daily log returns for the S&P 500 index from January 1, 2000 to December 30, 2021. The data were obtained from WRDS and the maximum likelihood estimates are presented in Table 2 along with their standard errors (in parentheses).
Table 2: Heston-Nandi GARCH Estimation Results for Daily Log-returns of SPX
[…]
1.9781 1.15 ×$\times 10^{-14}$ 0.7593 5.67×$\times 10^{-6}$ 185.5 17,963
(0.3166) (0.85×$\times 10^{-14}$) (0.0190) (1.18×$\times 10^{-6}$) (23.5)
Note: Maximum likelihood estimates of the Heston-Nandi GARCH model based on daily S&P 500 returns with robust standard errors in parentheses. Sample period: January 1, 2000 to December 30, 2021.
We can now illustrate Theorem 1 by computing the fractional absolute moments of cumulative returns, $R_{T,H}$ using the MGF we derived in Proposition 1. Figure 3 presents absolute moments, $\mathbb{E}|R_{T,H}|^{r}$, for a range of $r\in(-1,4]$ and $H=21$ (1 month), $H=63$ (3 month), and $H=126$ (6 months). The moments based on the new CMGF method are shown with colored lines, and these agree with the dashed lines, that are based on the average over 1,000,000 simulations. As for the NIG distribution, the CMGF is more accurate and more than 100 times faster than simulating 1,000,000 random and taking their average.
Figure 3: The conditional moments of absolute cumulative returns, $\mathbb{E}_{T}|R_{T,H}|^{r}$, plotted against $r$, for horizons: one month ($H=21$), three months ($H=63$), and six months ($H=126$). The initial value of $h_{T+1}$, is set as $\mathbb{E}(h_{t})$.
We can also illustrate Theorem 3 in this application, by computing the conditional integer moments, $\mathbb{E}_{T}[(R_{T,H})^{k}]$, $k\in\{1,2,3,4\}$. From these moments, we compute the conditional mean, $\mu_{T,H}$, standard deviation, $\sigma_{T,H}$, skewness, and kurtosis, where $\mu_{T,H}=\mathbb{E}_{T}[R_{T,H}]$, $\sigma_{T,H}^{2}=\mathbb{E}_{T}[R_{T,H}^{2}]-\mu_{T,H}^{2},$
$\operatorname{Skew}_{T,H}$ $=$ $\frac{\mathbb{E}_{T}[R_{T+H}^{3}]-\mu_{T,H}^{3}}{\sigma_{T,H}^{3}}-3\frac{\mu_{t,H}}{\sigma_{T,H}},$
$\operatorname{Kurt}_{T,H}$ $=$ $\frac{\mathbb{E}_{T}[R_{T+H}^{4}]-\mu_{T,H}^{4}}{\sigma_{T,H}^{4}}-4\frac{\mu_{T,H}}{\sigma_{T,H}}\operatorname{Skew}_{T,H}-6\frac{\mu_{T,H}^{2}}{\sigma_{T,H}^{2}}.$
Figure 4 plots these quantities against $H$ (solid lines) along with simulated quantities based on one million Monte Carlo simulations (dashed lines), where the initial value of $h_{T}$ is set to be the unconditional mean, $\mathbb{E}(h_{t})$. We find that the new expressions are in agreement with the simulated quantities.
Figure 4: Conditional moment quantities for cumulative returns, $R_{T,H}$ in the HNG model, plotted against $H$ that ranges from raging from one day to six months. The initial value of $h_{T+1}$ is set to be the unconditional mean, $\mathbb{E}(h_{t})$.
3.2 Moments in Autoregressive Gamma Model
In this application, $X_{t}$ represents the daily realized variance, which is computed from the intraday transaction data. We follow Corsi et al., (2013) and Majewski et al., (2015) and adopt a Heterogeneous Autoregressive Gamma (HARG) model for $X_{t}$. The HARG model is based on the Autoregressive Gamma (ARG) model by Gourieroux and Jasiak, (2006), and both employ a non-central Gamma distribution for the conditional density. [^6] The transition density that is implied by the CIR model, see Cox et al., (1985), is a non-central gamma density, and Engle and Gallo, (2006) employed the standard Gamma distribution as the conditional distribution in the MEM model. What sets the two apart, is that the HARG employs a long-memory structure for the location parameter. The conditional MGF conveniently has an affine form, which is practical for evaluating the integral expressions for the moments.
The HARG model has
$X_{t}|\mathcal{F}_{t-1}\sim f(x|\delta,\eta,\theta_{t})=\exp\left(-\frac{x}{\eta}-\theta_{t}\right)\left(\sum_{k=0}^{\infty}\frac{x^{\delta+k-1}}{\eta^{\delta+k}\Gamma(\delta+k)}\frac{\theta_{t}}{k!}\right),\quad x>0,$
which is the density for a non-central Gamma distribution with shape parameter $\delta>0$, scale parameter $\eta>0$, and location parameter $\theta_{t}>0$. The location parameter is time-varying, and given by
$\theta_{t}=\beta_{d}X_{t}^{(d)}+\beta_{w}X_{t}^{(w)}+\beta_{m}X_{t}^{(m)},$ (9)
where the variables, $X_{t}^{(d)}=X_{t-1}$, $X_{t}^{(w)}=\frac{1}{4}\sum_{i=2}^{5}X_{t-i}$, and $X_{t}^{(m)}=\frac{1}{17}\sum_{i=6}^{22}X_{t-i}$ represent lagged daily, “weekly”, and “monthly” averages. Observe that $\theta_{t}$ is $\mathcal{F}_{t-1}$-measurable. The HARG model implies a restricted AR(22) structure,
$\mathbb{E}_{t-1}[X_{t}]=\eta\delta+\eta\theta_{t}=\mu+\sum_{j=1}^{22}\phi_{j}X_{t-j},$ (10)
where $\mu=\eta\delta$ , $\phi_{1}=\eta\beta_{d}$, $\phi_{2}=\cdots=\phi_{5}=\tfrac{\eta}{4}\beta_{w}$, and $\phi_{6}=\cdots=\phi_{22}=\tfrac{\eta}{17}\beta_{m}$, as in the HAR model by Corsi, (2009).
The estimated parameters are reported in Table 3 along with their standard errors, where $\phi_{d}=\eta\beta_{d}=\phi_{1}$, $\phi_{w}=\eta\beta_{w}=\sum_{j=2}^{5}\phi_{j}$, and $\phi_{m}=\eta\beta_{m}=\sum_{j=6}^{22}\phi_{j}$, such that $\phi_{d}+\phi_{w}+\phi_{m}=\sum_{j=1}^{22}\phi_{j}$ is a measure of persistence.
The conditional MGF is conveniently given by
$M_{X_{T+1}|T}(z)=\exp\left(\frac{\eta z}{1-\eta z}\theta_{T+1}-\delta\log(1-\eta z)\right),\qquad\text{for}\quad\operatorname{Re}(z)<\eta^{-1}.$ (11)
From the estimated model we compute the term structure of conditional moments, $\mathbb{E}_{T}[X_{T+H}^{r}]$. The first conditional moment, $r=1$, is given directly from (10), and we will apply Theorem 2 to obtain other moments. The realized variance measures the second moment of returns, such that $r=\frac{1}{2}$ corresponds to the conditional volatility (standard deviation of returns). Similarly, $r=\frac{3}{2}$ and $r=2$ measures of the conditional skewness (of the absolute value) and the conditional kurtosis, respectively. The inverse of volatility corresponds to the negative moment, $r=-\frac{1}{2}$. This moment is interesting because it is used for Shape ratio forecasting and asset allocation, where the inverse of covariance matrix is employed. We compute these four moments by Theorem 2. The analytical form of the conditional MGF of $X_{T}$ is given in Proposition 2.
Proposition 2.
Suppose $X_{t}$ follows a HARG(p) process. Then the conditional MGF for $X_{T+H}$, given $\mathcal{F}_{T}$, is given by
$M_{X_{T+H}|T}(z)\equiv\mathbb{E}_{T}\left(e^{zX_{t+H}}\right)=\exp\left(A(H,z)+\sum_{j=1}^{p}B_{j}(H,z)X_{T+1-j}\right),$
where $A(H,z)$ and $B_{j}(H,z)$, are given from the initial values,
$A(1,z)=-\delta\log(1-\eta z)\quad\text{and}\quad B_{j}\left(1,z\right)=\frac{z}{1-\eta z}\phi_{i},\quad j=1,\ldots,p,$
and the recursions,
$A(h+1,z)$ $=$ $A(h,z)-\delta\log[1-\eta B_{1}(h,z)],$
$B_{j}\left(h+1,z\right)$ $=$ $\tfrac{B_{1}(h,z)}{1-\eta B_{1}(h,z)}\phi_{j}+1_{\{j<p\}}B_{j+1}(h,z),\qquad j=1,\ldots,p.$
Note that $M_{X_{T+H}|T}(z)$ is well defined for $z\in\{\zeta\in\mathbb{C}:B_{1}(h,\operatorname{Re}(\zeta))<\eta^{-1}\text{ forall }h\leq H\}$ and that $p=22$ in this model.
Table 3: HARG Estimation Results for Daily Realized Variance
[…]
0.4896 0.2789 0.0357 0.0053 0.9644 17,298
(0.0279) (0.0333) (0.0120) (0.0003) (0.0226)
Note: Maximum likelihood estimates for the HARG model with robust standard errors in parentheses. We report $\tilde{\beta}_{j}=\eta\beta_{j}$ for $j=\{d,w,m\}$ because they (unlike $\beta_{j}$) can be interpreted as the AR coefficients.
We adopt the realized (Parzen) kernel estimator, $\operatorname{RK}_{t}$ by Barndorff-Nielsen et al., (2008), as our realized measure of the variance. The daily $\operatorname{RK}_{t}$ is estimated for the S&P 500 index over the sample period from January 1, 2000 to November 30, 2021 (5,490 trading days). [^7] The data of realized kernel was obtained from the Realized Library at the Oxford-Man Institute, which was discontinued in 2022. The realized measures are annualized by the scaling, $X_{t}=252\times\operatorname{RK}_{t}$. Table 3 presents the maximum-likelihood estimates for the HARG models along with their corresponding standard errors (in parentheses).
Figure 5: The conditional $r$-th moment, $\mathbb{E}_{T}[(X_{T+H})^{r}]$, in HARG model for $r\in\{-\frac{1}{2},\frac{1}{2},\frac{3}{2},2\}$ and $H$ ranging from one day to six months. There is agreement between the CMGF moments (solid lines) and the simulated moments based on $N=10^{6}$ simulations (dashed line). The data generating process is the HARG model with parameter values set to the estimates in Table 3. Conditional moments are compute for $X_{T}=X_{T-1}=\ldots=\frac{1}{10}\mathbb{E}[X_{t}]$, where $\mathbb{E}\left(X_{t}\right)=\eta\delta/\left(1-\eta\beta_{d}-\eta\beta_{w}-\eta\beta_{m}\right)$.
Under Proposition 2, we compute the term structure of moments $\mathbb{E}_{t}\left(X_{T+H}^{r}\right)$ for $H=1,\ldots,180$ and $r\in\{-\frac{1}{2},\frac{1}{2},\frac{3}{2},2\}$. Figure 5 plots the term structure of these four moments when the initial value of lagged $X_{t}$ components are all set as $\tfrac{1}{10}\mathbb{E}[X_{t}]$, where $\mathbb{E}[X_{t}]=\eta\delta/\left(1-\eta\beta_{d}-\eta\beta_{w}-\eta\beta_{m}\right).$ We include both the simulated value (solid line) from one million Monte-Carlo simulations and the numerical value (dashed line) from Proposition 2. We can find the numerical values fit the simulated values very well.
3.3 Moments in Autoregressive Poisson Model
The autoregressive Poisson (ARP) model is given by
$\Pr(Y_{t}=y|\mathcal{F}_{t-1})=\frac{\lambda_{t}^{y}}{y!}e^{-\lambda_{t}},\qquad y=0,1,\ldots$ (12)
where the dynamic intensity parameter evolves according to
$\lambda_{t+1}=\omega+\beta\lambda_{t}+\alpha Y_{t},\qquad t=1,2,\ldots,$ (13)
such that $\lambda_{t+1}\in\mathcal{F}_{t}$. The average count over the next $H$ periods is denoted,
$\bar{Y}_{T,H}=\frac{1}{H}\sum_{h=1}^{H}Y_{T+h},$
and we seek the conditional moments of $\bar{Y}_{T,H}$ given $\mathcal{F}_{T}$. This could in principle be computed from the conditional distributions of $(Y_{T+1},\ldots,Y_{T+H})$, which can be inferred from the ARP model. However, there is substantial combinatorial complexity involved with this, and the complexity increases rapidly with $H$. The new method makes it simpler to compute moments, especially if $H$ is large.
Table 4: ARP Model Estimation Results for Count data of CBOE VIX Jumps
[…]
0.1548 0.7473 0.2043 3.2024 -9722
(0.0386) (0.0349) (0.0243)
Note: ML estimation results for the ARP model with robust standard errors in parentheses.
For $H=1$, the conditional MGF is
$M_{Y_{T+1}|T}(z)=\exp\left(\lambda_{T+1}\left(e^{z}-1\right)\right).$ (14)
More generally, the analytical form of the conditional MGF of $\bar{Y}_{T,H}$ is given in the following Proposition 3.
Proposition 3.
Let $Y_{t}$ be given by (12) and (13). Then
$M_{\bar{Y}_{T,H}|T}(z)=\exp\left(A(H,z)+B(H,z)\lambda_{T+1}\right),$
where $A(H,z)$ and $B(H,z)$ are given from
$A(h+1,z)$ $=A(h,z)+\omega B(h,z),$
$B(h+1,z)$ $=\beta B(h,z)+(e^{z/H+\alpha B(h,z)}-1),$
with initial value $A(1,z)=0$, and $B(1,z)=e^{z/H}-1$.
We estimate an ARP model for daily volatility jumps. Let $Y_{t}$ be the number of daily volatility jumps, as defined by intraday jumps in CBOE VIX index. We obtain high-frequency VIX data from Tick-data for the sample period from July 1, 2003 to December 30, 2021. We use the method by Andersen et al., (2010) to identity the number of daily volatility jumps. [^8] A similar approach was used in Alitab et al., (2020) to determine the number of jumps in S&P 500 index. We estimated the ARP model with maximum likelihood. The estimate parameters are presented in Table 3 along with their standard errors (in parentheses). The average jump intensity is about 3.2 jumps per day and $\hat{\pi}=\hat{\alpha}+\hat{\beta}=0.9517$ shows that the jump intensity is persistent.
Figure 6: Conditional moments, $\mathbb{E}_{t}[(\bar{Y}_{T,H})^{r}]$ in ARG model, where of $\bar{Y}_{T,H}=\frac{1}{H}\sum_{h=1}^{H}Y_{T+h}$. We present the four moments, $r\in\{\frac{1}{2},1,\frac{3}{2},2\}$, for $H$ ranging from 1 day to six months. We include both the simulated value (solid line) from one million Monte-Carlo simulations and the numerical value (dashed line) from Proposition 2. Model parameters are taken from Table 4. The initial value of $\lambda_{T+1}$ is set as $\frac{1}{10}\mathbb{E}\left(\lambda_{t}\right)$, where $\mathbb{E}\left(\lambda_{t}\right)=\omega/\left(1-\beta-\alpha\right)$. The horizontal axis indicates the calendar days.
Under Proposition 3, we compute the term structure of moments $\mathbb{E}_{T}[(\bar{Y}_{T,H})^{r}]$ for $r\in\{\frac{1}{2},1,\frac{3}{2},2\}$ and $H$ ranging from 1 day to six months. Figure 6 plots the term structure of these four moments when the initial value of $\lambda_{t+1}$ set as $\frac{1}{10}\mathbb{E}\left(\lambda_{t}\right)$. We include both the simulated value (solid line) from one million Monte-Carlo simulations and the numerical value (dashed line) from Proposition 3. We can find the numerical values fit the simulated values very well.
4 Conclusion
In this paper, we introduced a novel method for computing moments, including fractional moments, of random variables using their moment-generating function (MGF). A key advantage of our approach is that it avoids the need for MGF derivatives, which can be computationally challenging or unavailable in many models. We provided new integral expressions for fractional moments, fractional absolute moments, and central moments that extend the applicability of moment computation that is grounded in the MGF. The CMGF method leverages a complex extension of the MGF and is flexible enough to handle non-integer and complex moments.
The CMGF method may be valuable in structural models where moments play an important role. Moments and conditional moments are also central to inference methods. For instance, the generalized method of moments (GMM) by Hansen, (1982) requires the mapping from parameter to moments to be known. By offering solutions where other analytical methods fall short, the CMGF method can broaden the applicability of GMM to cover some problems that currently require simulated method of moments (SMM), see McFadden, (1989) and Duffie and Singleton, (1993).
We found the new method to be very fast and highly accurate for computing moments of the normal-inverse Gaussian distribution. Moreover, the CMGF method is especially useful in dynamic models where the MGF is known but derivatives are difficult to obtain, as demonstrated by the three applications in Section 3, where we computed moments of cumulative returns in a Heston-Nandi GARCH model, moments of realized volatilities in a Heterogeneous Autoregressive Gamma model, and moments of number of volatility jumps in an Autoregressive Poisson model.
Future research could explore further extensions of this method to other models with closed-form MGFs. An interesting extension is to apply the CMGF method to multivariate distributions, allowing for the computation of moments in multivariate distributions, including cross-moments that capture dependencies, beyond the simplest cross-moments, $\mathbb{E}[X_{1}X_{2}]$, discussed in Section 2.
Appendix A Appendix of Proofs
Some notation: We write +=$\mathbb{C}^{+}=\{z\in\mathbb{C}:\mathrm{Re}[z]>0\}$ to denote the complex numbers with positive real part.
In the proof of Theorem 1, we will need the following result (the Laplace transformation of $x^{r}$).
Lemma A.1.
Let $r+1\in\mathbb{C}^{+}$ and $z\in\mathbb{C}^{+}$. Then
$\int_{0}^{\infty}x^{r}e^{-zx}\mathrm{d}x$ $=\frac{\Gamma(r+1)}{z^{r+1}}.$ (A.1)
Proof.
Consider the case where $z$ is real and positive, i.e. $z=s>0$. From the definition of the gamma function and the substitution, $y\mapsto sx$, we have $\Gamma(r+1)=\int_{0}^{\infty}y^{r}e^{-y}\mathrm{d}y=\int_{0}^{\infty}s^{r}x^{r}e^{-sx}s\mathrm{d}x$, which proves the identity for any $z\in\mathbb{R}^{+}$. Both sides of (A.1) are holomorphic functions for $z\in\mathbb{C}^{+}$, and since they agree on $z\in\mathbb{R}^{+}$, it follows by the identity theorem that they agree for all $z\in\mathbb{C}^{+}$. [^9] We thank Christian Berg for suggesting this simplification of the proof. A different, direct proof for all $z\in\mathbb{C}^{+}$, is presented in the Supplementary Material. ∎
Lemma A.2.
For $x<0$, $\operatorname{Re}(z)>0$, and $\operatorname{Re}(r)>-1$ we have
$\int_{-\infty}^{+\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}t=0.$
For $x=0$, the identity holds for $\operatorname{Re}(r)>0$.
Proof.
We first consider the case $x=0$. The result follows directly from,
$\int_{-\infty}^{+\infty}\frac{e^{z\cdot 0}}{z^{r+1}}\mathrm{d}t=\int_{-\infty}^{+\infty}\frac{1}{(s+it)^{r+1}}\mathrm{d}t=\left.\frac{i}{r}\frac{1}{(s+it)^{r}}\right|_{-\infty}^{+\infty},$
because the last term is zero for ${\rm Re}\left(r\right)>0$.
Next, for $x<0$ we have
$\int_{-\infty}^{+\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}t$ $=\int_{-\infty}^{+\infty}\frac{e^{\left(s+it\right)x}}{\left(s+it\right)^{r+1}}\mathrm{d}t$
$=-\int_{+\infty}^{-\infty}\frac{e^{\left(s-i\tau\right)x}}{\left(s-i\tau\right)^{r+1}}\mathrm{d}\tau,\quad\tau=-t$
$=\int_{-\infty}^{+\infty}\frac{e^{\left(s-i\tau\right)x}}{\left(s-i\tau\right)^{r+1}}\mathrm{d}\tau$
$=e^{sx}\int_{-\infty}^{+\infty}\frac{e^{-i\tau x}}{\left(s-i\tau\right)^{r+1}}\mathrm{d}\tau$
$=\lim_{R\rightarrow\infty}e^{sx}\int_{-R}^{+R}f(\tau)\mathrm{d}\tau,$
where $f(\tau)=\frac{e^{-i\tau x}}{\left(s-i\tau\right)^{r+1}}$. The seemingly odd change of variables, $\tau=-t$, serves two important purposes that we explain below.
We will use a closed contour integral in the complex plane with two part. The first part is the line segment along the real axis, from $-R$ to $R$, and the second part is the semicircular arc, $C_{R}=\{Re^{i\theta},\theta\in[0,\pi]\}$. The integral over the line segment, $\int_{-R}^{+R}f(\tau)\mathrm{d}\tau$, is the object of interest, whereas the semicircle defines an auxiliary integral that is used to close the contour, such that we can apply the Cauchy Integral Theorem. The first reason for the change of variable, $\tau=-t$, is to place the pole of $f(\zeta)$ outside the closed contour for all $R>0$ (the pole is at $\zeta=-is$), such that $0=\int_{-R}^{R}f(\zeta)\mathrm{d}\zeta+\int_{C_{R}}f(\zeta)\mathrm{d}\zeta$ by the Cauchy Integral Theorem. For the integral along the arc, we have the expression,
$\int_{C_{R}}f(\zeta)\mathrm{d}\zeta=\int_{C_{R}}e^{-i\zeta x}g(\zeta)\mathrm{d}\zeta=\int_{C_{R}}e^{ia\zeta}g(\zeta)\mathrm{d}\zeta,$
where $a=-x>0$ and $g(\zeta)=\frac{1}{(s-i\zeta)^{r+1}}$ is continuous and satisfies $\lim_{|\zeta|\rightarrow\infty}g(\zeta)=0$ for all $r+1\in\mathbb{C}^{+}$. The second reason for the change of variable is that it enables us to apply Jordan’s lemma, which requires $a>0$ in the expression above. By Jordan’s lemma we have that $\lim_{R\rightarrow\infty}\int_{C_{R}}e^{-ix\zeta}g(\zeta)\mathrm{d}\zeta=0$, which implies that
$\int_{-\infty}^{+\infty}f(\tau)\mathrm{d}\tau=\lim_{R\rightarrow\infty}\int_{-R}^{+R}f(\tau)\mathrm{d}\tau=-\lim_{R\rightarrow\infty}\int_{C_{R}}f(\zeta)\mathrm{d}\zeta=0.$
This completes the proof. ∎
Proof of Lemma 1. Consider first the case $x\geq 0$. Since $\mathrm{Re}[z]=s>0$, we have the following identity from Lemma A.1,
$\int_{0}^{\infty}x^{r}e^{-zx}\mathrm{d}x$ $=$ $\frac{\Gamma(r+1)}{z^{r+1}},\qquad\text{for all}\quad\mathrm{Re}[r]>-1.$
This is the Laplace transform of $x^{r}$ and the inverse Laplace transform is
$x^{r}=\frac{1}{2\pi i}\int_{s-i\infty}^{s+i\infty}\frac{\Gamma(r+1)}{z^{r+1}}e^{zx}\mathrm{d}z,\quad x>0,$
provided the integral is well defined. Using $x>0$, we can rewrite the integral as
$|x|^{r}=x^{r}=\frac{\Gamma(r+1)}{2\pi i}\int_{s-i\infty}^{s+i\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}z=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}t,$ (A.2)
which is well-defined for all $\mathrm{Re}[r]>-1$ (because $x>0$). For $x=0$ the expression is well-defined for all $\mathrm{Re}[r]>0$. This completes the proof of the third expression in Lemma 1. We proceed with
$|x|^{r}$ $=x^{r}=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}t+\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-zx}}{z^{r+1}}\mathrm{d}t$
$=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}+e^{-zx}}{z^{r+1}}\mathrm{d}t,\qquad\qquad\quad x\geq 0,$ (A.3)
where the second term is zero by Lemma A.2.
For the case $x<0$, we have the following identity directly from (A.2),
$|x|^{r}=\left(-x\right)^{r}$ $=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-zx}}{z^{r+1}}\mathrm{d}t,\qquad x<0.$
Adding $\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}t$ to the right-hand side, which equals zero Lemma A.2, we arrive at
$|x|^{r}=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-zx}+e^{zx}}{z^{r+1}}\mathrm{d}t,\qquad\text{for }x<0.$ (A.4)
Combining (A.3) and (A.4) proves the first expression in Lemma 1.
From (A.2), we also have
$x^{k}$ $=\qquad\quad\frac{\Gamma(k+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{+zx}}{z^{k+1}}\mathrm{d}t,\quad x>0,$
$x^{k}$ $=\left(-1\right)^{k}\frac{\Gamma(k+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-zx}}{z^{k+1}}\mathrm{d}t,\quad x<0.$
This, and $\Gamma(k+1)=k!$ for $k\in\mathbb{N}$, proves the second expression in Lemma 1. This completes the proof. $\square$
Proof of Theorem 1. The proof is the same as in the proof the Theorem 2 by using the first formula in Lemma 1. $\square$
Proof of Theorem 2. We prove the result for discrete and continuous distributions separately. Consider first a non-negative discrete random variable, $X$, with support $\mathcal{X}$. Rewriting $\mathbb{E}[X^{r}]=\sum_{x\in\mathcal{X}}x^{r}p(x)$ using the third formula in Lemma 1 yields
$\mathbb{E}[X^{r}]$ $=\frac{\Gamma(r+1)}{2\pi}\sum_{x\in\mathcal{X}}\left[\int_{-\infty}^{\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}t\right]p(x)$
$=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{\infty}\sum_{x\in\mathcal{X}}\frac{e^{zx}}{z^{r+1}}p(x)\mathrm{d}t$
$=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{\infty}\frac{M_{X}\left(z\right)}{z^{r+1}}\mathrm{d}t,$
for $r+1\in\mathbb{C}^{+}$ if $\Pr(X=0)=0$ and $r\in\mathbb{C}^{+}$ if $\Pr(X=0)>0$. For a continuous distribution, the same line of arguments gives us,
$\mathbb{E}[X^{r}]=\tfrac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{\infty}\left[\int_{-\infty}^{\infty}\tfrac{e^{zx}}{z^{r+1}}\mathrm{d}t\right]f(x)\mathrm{d}x=\tfrac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{\infty}\tfrac{M_{X}\left(u\right)}{z^{r+1}}\mathrm{d}t.$
This completes the proof. $\square$
Proof of Theorem 3. The proof is the same as in the proof the Theorem 1 by using the second formula in Lemma 1. $\square$
Appendix B Result for the Dynamic Models
B.1 MGF for ARG model
Proof of Proposition 1. First, the dynamic process of $\theta_{t}$ can be rewritten as
$\theta_{T}=\sum_{j=1}^{p}\beta_{j}X_{T-j},\quad\beta_{i}=\begin{cases}\beta_{d}&j=1\\ \beta_{w}/4&2\leq j\leq 5\\ \beta_{m}/17&6\leq j\leq 22\end{cases}$
where $p=22$. For $T=T+H$, assume that the conditional MGF for $X_{T+H}$ is given by
$M_{T,H}\left(z\right)=\mathbb{E}_{T}\left(e^{zX_{T+H}}\right)=\exp\left(A(H,z)+\sum_{j=1}^{p}B_{j}\left(H,z\right)X_{T+1-j}\right).$
For the case, $H=1$, we can use the MGF (11),
$\mathbb{E}_{T}\left(e^{zX_{T+1}}\right)$ $=\exp\left(\delta\log\tfrac{1}{1-\eta z}+\tfrac{\eta z}{1-\eta z}\theta_{T+1}\right)=\exp\left(A(1,z)+\sum_{j=1}^{p}B_{j}(1,z)X_{T+1-j}\right),$
for $s<\eta^{-1}$ where
$A(1,z)=-\delta\log(1-\eta z),\quad B_{j}\left(1,z\right)=\frac{\eta z}{1-\eta z}\beta_{j},\quad j=1,\ldots,p.$
The recursions (from $h$ to $h+1$) follow from
$\mathbb{E}_{T}\left(e^{zx_{T+h+1}}\right)$ $=\mathbb{E}_{T}\left[\mathbb{E}_{T+1}\left(e^{zX_{T+h+1}}\right)\right]$
$=\mathbb{E}_{T}\left[\exp\left(A(h,z)+\sum_{j=1}^{p}B_{j}(h,z)X_{T+2-j}\right)\right]$
$=\exp\left(A(h,z)+\sum_{j=2}^{p}B_{j}(h,z)X_{T+2-j}\right)\mathbb{E}_{T}\left[\exp\left(B_{1}(h,z)X_{T+1}\right)\right]$
$=\exp\left(A(h,z)+\sum_{j=1}^{p-1}B_{j+1}(h,z)X_{T+1-j}\right)$
×$\quad\times\exp\left(-\delta\log\left(1-\eta B_{1}(h,z)\right)+\frac{\eta B_{1}(h,z)}{1-\eta B_{1}(h,z)}\sum_{j=1}^{p}\beta_{j}X_{T+1-j}\right)$
$=\exp\left(A(h+1,z)+\sum_{j=1}^{p}B_{j}(h+1,z)x_{T+1-j}\right),$
where
$A(h+1,z)=A(h,z)-\delta\log(1-\eta B_{1}\left(h,z\right))$
and
$B_{j}\left(h+1,z\right)=\begin{cases}B_{j+1}(h,z)+\frac{\eta B_{1}(h,z)}{1-\eta B_{1}(h,z)}\beta_{j}&1\leq j<p\\ \frac{\eta B_{1}(h,z)}{1-\eta B_{1}(h,z)}\beta_{j}&j=p\end{cases}$
Note that we need the $s$ to be such that
$B_{1}\left(h,s\right)<\eta^{-1}\quad{\rm for}\quad\forall\ h\leq T-t.$
That completes the proof of Proposition 2. $\square$
Conditional MGF for Autoregressive Poisson
Proof of Proposition 2. The conditional MGF for a Poisson distributed random variable, $Y_{t}$, with intensity $\lambda_{t}$, is $\mathbb{E}_{t-1}\left(e^{zY_{t}}\right)=\exp\left(\lambda_{t}\left(e^{z}-1\right)\right)$. So for the ARP process,
$Y_{t}|\mathcal{F}_{t-1}\sim\operatorname{Poisson}(\lambda_{t}),\qquad\text{with}\quad\lambda_{t}=\omega+\beta\lambda_{t-1}+\alpha n_{t-1}.$
The average number over the next $H$ periods, is
$\bar{Y}_{T,H}=\frac{1}{H}S_{T,H},\quad\text{where}\quad S_{T,H}=\sum_{h=1}^{H}Y_{T+h}.$
We will show, by induction, that $M_{S_{T,H}}(z)$ has an affine structure. For $H=1$, we have $S_{T,1}=Y_{T+1}$, such that $M_{S_{T,1}}(z)=\mathbb{E}_{t}[e^{zY_{T+1}}]=\exp[\lambda_{T+1}(e^{z}-1)]$, which can be expressed in the affined form,
$M_{S_{T,1}}(z)=\exp\left(\tilde{A}(1,z)+\tilde{B}(1,z)\lambda_{T+1}\right),\quad\text{with }\tilde{A}(1,z)\equiv 0\text{ and }\tilde{B}(1,z)\equiv e^{z}-1.$
Next, suppose that $M_{S_{T,h}}(z)=\exp(\tilde{A}(h,z)+\tilde{B}(h,z)\lambda_{T+1})$ and consider
$M_{S_{T,h+1}}(z)$ $=$ $\mathbb{E}_{T}[\exp(zS_{T,h+1})]=\mathbb{E}_{T}[\exp(zY_{T+1}+zS_{T+1,h})]$
$=$ $\mathbb{E}_{T}[\mathbb{E}_{T+1}\{\exp(zY_{T+1}+zS_{T+1,h})\}]$
$=$ $\mathbb{E}_{T}[\exp\{zY_{T+1}+\tilde{A}(h,z)+\tilde{B}(h,z)\lambda_{T+2}\}]$
$=$ $\mathbb{E}_{T}[\exp\{zY_{T+1}+\tilde{A}(h,z)+\tilde{B}(h,z)(\omega+\beta\lambda_{T+1}+\alpha Y_{T+1})\}]$
$=$ $\exp\{\tilde{A}(h,z)+\tilde{B}(h,z)(\omega+\beta\lambda_{T+1})\}\mathbb{E}_{T}[\exp\{(z+\alpha\tilde{B}(h,z))Y_{T+1})\}]$
$=$ $\exp\{\tilde{A}(h,z)+\tilde{B}(h,z)(\omega+\beta\lambda_{T+1})\}\exp\{\lambda_{T+1}(e^{z+\alpha\tilde{B}(h,z)}-1)\}$
$=$ $\exp\{\tilde{A}(h,z)+\tilde{B}(h,z)(\omega+\beta\lambda_{T+1})+\lambda_{T+1}(e^{z+\tilde{B}(h,z)\alpha}-1)\}$
$=$ $\exp\{[\underset{=\tilde{A}(h+1,z)}{\underbrace{\tilde{A}(h,z)+\omega\tilde{B}(h,z)}}]+[\underset{\tilde{B}(h+1,z)}{\underbrace{\beta\tilde{B}(h,z)+(e^{z+\tilde{B}(h,z)\alpha}-1)}}]\lambda_{T+1}\}.$
This proves that $M_{S_{T,h+1}}(z)$ has an affine structure and, by an induction argument, so does $M_{S_{T,h+2}}(z),\ldots,M_{S_{T,H}}(z)$. Next, $M_{\bar{Y}_{T,H}}(z)=M_{S_{T,H}}(z/H)$. So, if we set $A(h,z)\equiv\tilde{A}(h,z/H)$ and $B(h,z)=B(h,z/H)$, then $M_{\bar{Y}_{T,H}}(z)=\exp[A(H,z)+B(H,z)]\lambda_{T+1}$ has the affine structure, where $A(h+1,z)=A(k,z)+\omega B(k,z)$ and $B(h+1,z)=\beta B(h,z)+\left(e^{z/H+\alpha B(h,z)}-1\right)$, with initial values $A(1,z)=0$ and $B(1,z)=(e^{z/H}-1)$. $\square$
B.2 The MGF in Heston-Nandi GARCH Model
We first have
$\mathbb{E}_{t}\left[\exp\left(v_{1}r_{t+1}+v_{2}h_{t+2}\right)\right]=$ $\exp\left[v_{1}r+v_{2}\omega+\left[v_{1}\left(\lambda-\tfrac{1}{2}\right)+v_{2}\left(\beta+\alpha\gamma^{2}\right)\right]h_{t+1}\right]$
×$\times\mathbb{E}_{t}\left[v_{2}\alpha z_{t+1}^{2}+\left(v_{1}-2v_{2}\alpha\gamma\right)\sqrt{h_{t+1}}z_{t+1}\right]$
$=$ $\exp\left[v_{1}r+v_{2}\omega+\left[v_{1}\left(\lambda-\tfrac{1}{2}\right)+v_{2}\left(\beta+\alpha\gamma^{2}\right)\right]h_{t+1}\right.$
$\left.-\frac{1}{2}\log\left(1-2v_{2}\alpha\right)+\frac{\left(v_{1}-2v_{2}\alpha\gamma\right)^{2}}{2\left(1-2v_{2}\alpha\right)}h_{t+1}\right]$
$=$ $\exp\left[v_{1}r+v_{2}\omega-\tfrac{1}{2}\log\left(1-2v_{2}\alpha\right)\right.$
$\left.+\left(v_{1}\left(\lambda-\tfrac{1}{2}\right)+v_{2}\left(\beta+\alpha\gamma^{2}\right)+\frac{\left(v_{1}-2v_{2}\alpha\gamma\right)^{2}}{2\left(1-2v_{2}\alpha\right)}\right)h_{t+1}\right].$
Next, we show by induction that the conditional MGF has the form,
$\mathbb{E}_{t}\left(\exp\left(s\sum_{i=1}^{H}r_{t+i}\right)\right)=\exp\left(A(H,s)+B(H,s)h_{t+1}\right).$
For $h=1$ we have
$\mathbb{E}_{t}\left(\exp\left(sr_{t+1}\right)\right)$ $=\exp\left(sr+\left(s\left(\lambda-\tfrac{1}{2}\right)+\tfrac{s^{2}}{2}\right)h_{t+1}\right)=\exp\left(A(1,s)+B(1,s)h_{t+1}\right)$
such that $A(1,s)=sr$ and $B(1,s)=s\left(\lambda-\tfrac{1}{2}\right)+\frac{s^{2}}{2}$. Next, suppose that the MGF has an affine structure for $h$, then for $h+1$ we find
$\mathbb{E}_{t}\left(\exp\left(s\sum_{i=1}^{h+1}r_{t+i}\right)\right)$ $=\mathbb{E}_{t}\left[\mathbb{E}_{t+1}\left(\exp\left(s\sum_{i=1}^{h+1}r_{t+i}\right)\right)\right]$
$=\mathbb{E}_{t}\left[\exp\left(sR_{t+1}\right)\mathbb{E}_{t+1}\left(\exp\left(s\sum_{i=1}^{h}r_{t+1+i}\right)\right)\right]$
$=\mathbb{E}_{t}\left[\exp\left(sR_{t+1}+A(h,s)+B(h,s)h_{t+2}\right)\right]$
$=\exp\left(A(h,s)+sr+B(h,s)\omega-\frac{1}{2}\log\left(1-2B(h,s)\alpha\right)\right)$
×$\quad\times\exp\left(\left(s\left(\lambda-\frac{1}{2}\right)+B(h,s)\left(\beta+\alpha\gamma^{2}\right)+\frac{\left(s-2B(h,s)\alpha\gamma\right)^{2}}{2\left(1-2B(h,s)\alpha\right)}\right)h_{t+1}\right)$
$=\exp\left(A(h+1,s)+B(h+1,s)h_{t+1}\right),$
where
$A(h+1,s)$ $=A(h,s)+sr+B(h,s)\omega-\frac{1}{2}\log\left(1-2B(h,s)\alpha\right)$
$B(h+1,s)$ $=s\left(\lambda-\frac{1}{2}\right)+B(h,s)\left(\beta+\alpha\gamma^{2}\right)+\frac{\left(s-2B(h,s)\alpha\gamma\right)^{2}}{2\left(1-2B(h,s)\alpha\right)},$
for any sufficiently small $s=\operatorname{Re}(z)>0$, such that $B(h,s)<\frac{1}{2\alpha}$ for all $h=1,\ldots,H$, to ensure that $\log\left(1-2B(h,s)\alpha\right)\in\mathbb{R}$. This completes the proof. $\square$
Supplement
An Direct Proof of Lemma A.1.
Proof of Lemma A.1. By the substitution, $y=zx\in\mathbb{C}$, we have
$\lim_{K\rightarrow\infty}\int_{0}^{K}x^{r}e^{-zx}\mathrm{d}x=\frac{1}{z^{r+1}}\int_{0}^{sK+itK}y^{r}e^{-y}\mathrm{d}y.$
The function $f(y)\equiv y^{r}e^{-y}$, has a pole at $y=0$ if $r<0$, and is a multi-valued function when $r$ is a non-integer. We will adopt the principal value by restricting the angle of $y=\lambda e^{i\theta}$ to $\theta\in(-\pi,\pi]$, and construct a closed contour integral that avoids the pole at $y=0$. The contour is illustrated in Figure S.1.
Figure S.1: The closed contour for the integral, $\oint y^{r}e^{-y}\mathrm{d}y$.
The function, $f(y)$, is analytic over the closed contour for $\lambda>0$, (i.e. it does not include any singularity). Therefore, by the Cauchy Integral Theorem, we have
$0=\oint f(y)\mathrm{d}y$ $=\int_{AB}f(y)\mathrm{d}y+\int_{BC}f(y)\mathrm{d}y+\int_{CD}f(y)\mathrm{d}y+\int_{DA}f(y)\mathrm{d}y,$
where the integral we seek is $-\int_{CD}f(y)\mathrm{d}y$ as $\lambda\rightarrow 0$ and $K\rightarrow\infty$.
For the integral over the arc, DA, we have $y=\lambda e^{i\theta}=\lambda\left(\cos\theta+i\sin\theta\right)$, and $dy=i\lambda e^{i\theta}d\theta$, such that
$\left|I_{DA}\right|=\left|\int_{DA}y^{r}e^{-y}\mathrm{d}y\right|$ $=\left|\int_{\pi/2}^{0}\lambda^{r}e^{i\theta r}e^{-\lambda\left(\cos\theta+i\sin\theta\right)}i\lambda e^{i\theta}\mathrm{d}\theta\right|$
$\leq\left|\lambda^{r+1}\right|e^{|\lambda|}\left|\int_{\pi/2}^{0}e^{i\theta r}e^{-i\lambda\sin\theta}ie^{i\theta}\mathrm{d}\theta\right|$
$\leq\left|\lambda^{r+1}\right|e^{|\lambda|}\left|\int_{\pi/2}^{0}\mathrm{d}\theta\right|=\tfrac{\pi}{2}\lambda^{{\rm Re}\left(r\right)+1}e^{|\lambda|}\rightarrow 0,\quad\text{as }\lambda\rightarrow 0\text{ for }{\rm Re}\left(r\right)>-1.$
For the BC-integral, we have $y=sK+iuK$ and $\mathrm{d}y=iK\mathrm{d}u$, such that
$\left|I_{BC}\right|=\left|\int_{sK}^{sK+itK}y^{r}e^{-y}\mathrm{d}y\right|$ $=\left|iK\int_{0}^{t}\left(sK+iuK\right)^{r}e^{-\left(sK+iuK\right)}\mathrm{d}u\right|$
$=\left|iK^{r+1}e^{-sK}\int_{0}^{t}\left(s+iu\right)^{r}e^{-iuK}\mathrm{d}u\right|$
$\leq K^{{\rm Re}\left(r\right)+1}e^{-sK}\left|\int_{0}^{t}\left|\left(s+iu\right)^{r}\right|\mathrm{d}u\right|\rightarrow 0,\qquad\text{as }K\rightarrow\infty,$
because $e^{-sK}$ decays faster than $K^{r+1}$ increases. So, as $K\rightarrow\infty$ and $\lambda\rightarrow 0$, we have
$0=\lim_{K\rightarrow\infty}\lim_{\lambda\rightarrow 0}\left(\int_{AB}f(y)\mathrm{d}y+\int_{CD}f(y)\mathrm{d}y\right)=\lim_{K\rightarrow\infty}\left(\int_{0}^{sK}f(y)\mathrm{d}y+\int_{sK+itK}^{0}f(y)\mathrm{d}y\right).$
The integral we seek is therefore given by,
$\lim_{K\rightarrow\infty}\int_{0}^{sK+itK}f(y)\mathrm{d}y=\lim_{K\rightarrow\infty}\int_{0}^{sK}f(y)\mathrm{d}y=\Gamma(r+1).$
This completes the proof. $\square$
NIG Moments
The shape of the integrands provide some intuition for the advantage of the CMGF method in the evaluation of moments of the NIG distribution with $(\xi,\chi)=(\frac{1}{2},-\frac{1}{3})$. The CMGF integrand is non-negligible over a domain that is very similar for all values of $r$ and the range of the integrand is also similar across $r$. In fact, all CMGF integrands have the same maximum value at zero. For the integrand, $|x|^{r}f(x)$, there is far more variation across different moments, $r$, which may explain that this method is slower than the CMGF method.
Figure S.2: Integrands for some selected NIG moments.
Existing Methods for Fractional Moments
Existing expressions for positive fractional moments include,
$\mathbb{E}|X|^{r}=C_{K}\int_{0}^{\infty}u^{-(1+r)}\left[-{\rm Re}[\varphi_{X}(u)]+\sum_{k=0}^{K}\frac{u^{2k}}{(2k)!}\varphi_{X}^{(2k)}(0)\right]\mathrm{d}u,$
where $K=\lfloor\tfrac{r}{2}\rfloor$and $C_{K}$ is a positive constant, see Kawata, (1972, theorem 11.4.4), and
$\mathbb{E}|X|^{r}$ $=\tfrac{\lambda}{\Gamma(1-\lambda)}[\cos(\tfrac{\pi r}{2})]^{-1}\int_{0}^{\infty}\frac{\varphi_{X}^{(k)}(0)-\varphi_{X}^{(k)}(u)}{u^{1+\lambda}}\mathrm{d}u,$ (S.1)
where $k=\lfloor r\rfloor$ and $\lambda=r-k$, see Laue, (1980).
Closely relate to (S.1) is the following expression for non-negative random variables,
$\mathbb{E}[X^{r}]=(-1)^{k}\tfrac{\lambda}{\Gamma(1-\lambda)}\int_{0}^{\infty}\frac{M_{X}^{(k)}(0)-M_{X}^{(k)}(-u)}{u^{1+\lambda}}\mathrm{d}u,\qquad(X\geq 0),$
where $k=\lfloor r\rfloor$ and $\lambda=r-k$, see Schürger, (2002), and another expression for non-negative variables is
$\mathbb{E}[X^{r}]$ $=\frac{1}{\Gamma(\tilde{\lambda})}\int_{0}^{\infty}u^{\tilde{\lambda}-1}M_{X}^{(\tilde{k})}(-u)\mathrm{d}u,\qquad(X\geq 0),$
where $\tilde{k}=\lceil r\rceil>0$ and $\tilde{\lambda}=\tilde{k}-r\in[0,1)$, see Cressie and Borkent, (1986). For strictly positive variables we have the following expression for negative fractional moments, $r<0$,
[…]
see Schürger, (2002, theorem 1.1).
Code for Implementation
Moments in HNG Model
% MATLAB (R2024a version)
function Y = SimulationMoments(N,H,r,model)
Omega = model(1);
Beta = model(2);
Alpha = model(3);
Gamma = model(4);
Lam = model(5);
rf = model(6);
h1 = model(7);
Z = randn(N,H);
h = zeros(N,H);
h(:,1) = h1;
for t=1:H-1
h(:,t+1)=Omega+Beta*h(:,t)+Alpha*(Z(:,t)-Gamma*sqrt(h(:,t))).^2;
end
Rmat = H*rf+(Lam-1/2)*sum(h,2)+sum(sqrt(h).*Z,2);
Y = mean(abs(Rmat).^r);
end
function Y=CMGF(H,r,model,Tol)
Y= gamma(r+1)/pi*integral(@(s) IntFunction(s,H,r,model),0,Inf,’AbsTol’,Tol,’ArrayValued’,false);
end
function fs = IntFunction(s,H,r,model)
Omega = model(1);
Beta = model(2);
Alpha = model(3);
Gamma = model(4);
Lam = model(5);
rf = model(6);
h1 = model(7);
uR = 5;
u = uR+s*complex(0,1);
v = -u;
A1 = u*rf;
B1 = u*(Lam-0.5)+u.^2/2;
A2 = v*rf;
B2 = v*(Lam-0.5)+v.^2/2;
for m = 1:H-1
A1 = A1+u*rf+B1*Omega-0.5*log(1-2*Alpha*B1);
A2 = A2+v*rf+B2*Omega-0.5*log(1-2*Alpha*B2);
B1 = u*(Lam-0.5)+B1*(Beta+Alpha*Gamma^2)+0.5*(u-2*B1*Alpha*Gamma).^2./( 1-2*Alpha*B1);
B2 = v*(Lam-0.5)+B2*(Beta+Alpha*Gamma^2)+0.5*(v-2*B2*Alpha*Gamma).^2./( 1-2*Alpha*B2);
end
fs = real(1./(u.^(r+1)).*(exp(A1+B1*hp1Q)+exp(A2+B2*hp1Q)));
end

### v4

@@
 Abstract
-We introduce a novel method for obtaining a wide variety of moments of a random variable with a well-defined moment-generating function (MGF).
+We introduce a novel method for obtaining a wide variety of moments of any random variable with a well-defined moment-generating function (MGF).
 We derive new expressions for fractional moments and fractional absolute moments, both central and non-central moments.
-The new moment expressions are relatively simple integrals that involve the MGF, but do not require its derivatives.
+The expressions are relatively simple integrals that involve the MGF, but do not require its derivatives.
 We label the new method CMGF because it uses a complex extension of the MGF and can be used to obtain complex moments.
@@
 1 Introduction
-Moments of random variables, including conditional moments, are central to much of economics, finance, and econometrics.
+Moments of random variables, including conditional moments, are central to many areas of hard and social sciences.
 In this paper, we introduce a new method for obtaining moments of random variables with a well-defined moment-generating function (MGF).
-The proposed method is general, computationally efficient, and can offer solutions to problems where standard methods are inadequate.
+The proposed method is general, computationally efficient, and can offer solutions to problems where existing methods are inadequate.
 It is well known that the $k$-th integer moment of a random variable $X$, with MGF, $M_{X}(s)=\mathbb{E}[e^{sX}]$, is given by $\mathbb{E}[X^{k}]=M_{X}^{(k)}(0)$, where $M_{X}^{(k)}(s)$ is the $k$-th derivative of $M_{X}(s)$.
-Non-integer moments and fractional absolute moments, $\mathbb{E}|X^{r}|$, $r\in\mathbb{R},$ have more complicated expressions, typically integrals that involve derivatives of the MGF or derivatives of the characteristic function (CF).
+Non-integer moments and fractional absolute moments, $\mathbb{E}|X|^{r}$, $r\in\mathbb{R},$ have more complicated expressions, typically integrals that involve derivatives of the MGF or derivatives of the characteristic function (CF).
 [^1] Kawata, (1972) provides an expression for $\mathbb{E}|X|^{r}$ that involves several derivatives of the CF and Laue, (1980) provides an expression based on fractional derivatives, see also Wolfe, (1975), Samko et al., (1993), Matsui and Pawlas, (2016), and Tomovski et al., (2022).
-For fractional moments of non-negative variables, similar expressions were derived in Cressie et al., (1981), Cressie and Borkent, (1986), Jones, (1987), and Schürger, (2002).
+For fractional moments of non-negative variables, similar expressions were derived in Cressie et al., (1981), Cressie and Borkent, (1986), Jones, (1987), Schürger, (2002), and Meng, (2005).
 In this paper, we derive novel integral expressions for computing fractional moments, including fractional absolute moments, and fractional central moments.
-Unlike existing expressions, the new expressions do not involve derivatives (of the MGF or the CF).
-Thus, the new expressions are particularly useful when derivatives are prohibitively difficult to obtain.
-This is often the case in dynamic models, where the conditional MGF is defined from recursive non-linear expressions.
-The new method is labelled CMGF because it relies on a complex extension of the MGF, and provides expressions for complex moments, $\mathbb{E}|X|^{r}$, $r\in\mathbb{C}$.
+The new expressions stand out by being applicable to all random variables with a MGF and can be used to compute moments of $Y=g(X)$, whenever $g(x)$ can be expressed as a linear combination of a finite number of fractional, integer, absolute, positive-part, and possibly centralized powers of $x$.
+The new expressions are particularly useful when derivatives are prohibitively difficult to obtain, as is often the case in dynamic models, where the conditional MGF is defined from recursive non-linear expressions.
+The new method is also useful in situations where the density is unavailable, as can be the case for compound distributions, exponential tilting, and some Lévy processes.
+The new method is labeled CMGF because it relies on a complex extension of the MGF, and provides expressions for complex moments, $\mathbb{E}|X|^{r}$, $r\in\mathbb{C}$ with $\operatorname{Re}(r)>-1$.
 Although complex moments are not commonly used in econometrics, this generalization can be included with minimal adaptation.
-[^2] Complex moments are commonly used other fields, including statistical mechanics, number theory, and quantum physics.
-In part, because complex moments are related to Mellin transforms.
-The new expressions can also be used for establishing new identities, such as the expression for the reciprocal Gamma function, which we infer from the absolute moment of a Gaussian random variable.
+[^2] Complex moments are commonly used other fields, including quantum physics, number theory, and statistical mechanics.
 The new CMGF expressions involve integrals, and while these cannot be evaluated analytically in most cases, they facilitate new ways to compute moments numerically, which are fast and accurate.
-We demonstrate this with the normal-inverse Gaussian distribution.
+We demonstrate this with the normal-inverse Gaussian (NIG) distribution.
+Existing expressions for absolute fractional moments of the NIG distribution are limited to $\mu$-centered moments, and the expressions are infinite series involving modified Bessel functions of the second kind.
+The CMGF integral expressions are simpler and do not involve special functions.
 We also show that the CMGF method offers fast and accurate moments in dynamic models, where we are not aware of good alternative methods.
 Specifically, we use the CMGF method to compute moments of cumulative returns in the Heston-Nandi GARCH (HNG) model, to compute moments of realized volatilities in the Heterogeneous Autoregressive Gamma (HARG) model, and to compute conditional moments in the Autoregressive Poisson Model.
+A key step in our proof is the identity, $\int_{-\infty}^{+\infty}\frac{e^{-itx}}{(s+it)^{r+1}}\mathrm{d}t=0$, that holds for $s>0$, $x>0$ and $\operatorname{Re}(r)>-1$.
+This identity simplifies expressions involving an inverse Laplace transform, such that the moment-generating function emerges.
+The identity, see Lemma 1, appears to be new, which can explain that these simple and general integral expressions were not discovered earlier.
+We are not the first to derive integral expressions that involve the MGF.
+Meng, (2005) derived integral expressions of $\mathbb{E}[X^{a}/Y^{b}]$, involving their joint MGF $M_{X,Y}(s_{1},s_{2})$ and its derivatives.
+The possibility of combining characteristic and moment-generating functions, $\mathbb{E}[e^{sY+itX}]$, is mentioned in the epilogue of Meng, (2005).
+For $X=Y$ this is an extension of the MGF (or the CF) to the complex plane, $M_{X}(s+it)$, which is at the core of our expression.
+An auxiliary result that arises from applying the integral expression to standard normal distribution, is a holomorphic integral representation of Gamma function and its reciprocal, which do not rely on analytic continuation, see Hansen and Tong, (2025).
 The remainder of this paper is organized as follows.
 We present the new moment expressions in Section 2.
-Theorem 1 has the expression for absolute fractional moments for general random variables, Theorem 2 has the expression for fractional moments of non-negative random variables, and Theorem 3 presents a new expression for integer moments.
-As an example, we apply CMGF to the normal-inverse Gaussian distribution.
+Theorem 1 has the expression for absolute fractional moments for general random variables, Theorem 2 presents positive-part moments, and Theorem 3 presents a new expression for integer moments.
+Corollary 1 has expressions for fractional moments of nonnegative (positive) random variables and Corollary 2 has expression for (odd integers) negative tail moments.
+We apply CMGF to the normal-inverse Gaussian distribution.
 Section 3 presents three applications of the new expressions in dynamic models, and Section 4 concludes.
 2 A New Method for Computing Moments (CMGF)
-We will use the following notation: $\mathbb{N}$ denotes the positive integers and $\mathbb{N}_{0}=\{0\}\cup\mathbb{N}$.
-We use $\mathcal{N}$ to denote a neighborhood about zero, for which $(-\delta,\delta)\subset\mathcal{N}$ for some $\delta>0$, and the positive part of $\mathcal{N}$ is denoted $\mathcal{N}_{+}\equiv\{s\in\mathcal{N}:s>0\}$.
-We use $z=s+it$ and $\zeta=\sigma+i\tau$ to represent complex variables, where $s=\operatorname{Re}(z)$ and $t=\operatorname{Im}(z)$, and similarly for $\zeta$.
-To avoid possible confusion, we use $\mathbb{E}[X^{k}]$ to denotes integer moments, $k\in\mathbb{N}$, whereas general moments (including fractional and complex moments) are represented with $\mathbb{E}[X^{r}]$, $r\in\mathbb{C}$.
-Our results are applicable to any random variable with a well-defined MGF, as characterized in the following assumption.
-Assumption 1.
-$X$ is a real random variable with
-$M_{X}(s)=\mathbb{E}[e^{sX}]<\infty,\qquad\text{for all}\quad s\in\mathcal{N},$
-where $\mathcal{N}$ is a neighborhood about zero.
-The MGF is typically used as a real function with a real argument, such that both $s$ and $M_{X}(s)$ take values in $\mathbb{R}.$ In this paper, we establish our results by means of complex arguments, $z\in\mathbb{C},$ such that $M_{X}(z)$ also takes values in $\mathbb{C}$.
+We use the following notation: $\mathbb{N}$ denotes the positive integers and $\mathbb{N}_{0}=\{0\}\cup\mathbb{N}$.
+We use $z=s+it\in\mathbb{C}$, where $s=\operatorname{Re}(z)>0$.
+We use $\mathbb{E}[X^{k}]$ to denote integer moments, $k\in\mathbb{N}_{0}$, whereas $\mathbb{E}[X^{r}]$, denotes general moments including fractional and complex moments, $r\in\mathbb{C}$.
+The MGF is typically introduced as a real function with a real argument, such that both $s$ and $M_{X}(s)$ take values in $\mathbb{R}.$ In this paper, we establish our results by means of complex arguments, $z\in\mathbb{C},$ such that $M_{X}(z)$ also takes values in $\mathbb{C}$.
+Interpretation and scope: For complex powers $z^{-(r+1)}$, we fix the principal branch of the complex logarithm (branch cut on $(-\infty,0]$) and interpret integrals as oscillatory (i.e., symmetric Cauchy principal values) whenever $1<\operatorname{Re}(r)\leq 0$.
+Assumptions on one-sided exponential moments of $X$ (i.e., $\mathbb{E}[e^{\pm sX}]<\infty$ for some $s>0$) are stated with each theorem.
+When $\operatorname{Re}(r)\in(-1,0)$, we additionally assume $\mathbb{E},\lvert X-\xi\rvert^{r}<\infty$ to justify exchanging expectation and integration.
 To gain some intuition for the complex extension of the MGF, we observe that
@@
 nests both the standard moment-generating function, $\psi(s,0)=M_{X}(s)$, and the characteristic function, $\psi(0,t)=\varphi_{X}(t)\equiv\mathbb{E}[e^{itX}]$, as special cases.
-The complex valued MGF, $M_{X}(z)$, is the bilateral Laplace–Stieltjes transform (apart from the sign of $z$), and for continuous random variables it is similar to the Laplace transform.
+The complex-valued MGF, $M_{X}(z)$, is the bilateral Laplace–Stieltjes transform (apart from the sign of $z$), and for continuous random variables it is similar to the Laplace transform.
 [^3] The Laplace transform of the density function, $f_{X}(x)$, is $\int_{0}^{\infty}e^{-zx}f_{X}(x)\mathrm{d}x$, and it differs from $M_{X}(z)$ in terms of the domain of integration and the sign of the exponent.
-Moreover, $M_{X}(z)$ is obviously also applicable to discrete distributions and distributions that are mixtures of continuous and discrete distributions.
-The results are based on the following identities.
+Moreover, $M_{X}(z)$ is applicable to discrete distributions and mixtures of continuous and discrete distributions.
+The following result is key to several simplifications.
 Lemma 1.
-Let $x\in\mathbb{R}$ and $z=s+it\in\mathbb{C}$, where $s>0$ is an arbitrary positive constant.
-$|x|^{r}=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}+e^{-zx}}{z^{r+1}}\mathrm{d}t,\qquad\operatorname{Re}(r)>-1,$
-for any $x\neq 0$, and this identity also holds for $x=0$ if $\operatorname{Re}(r)>0$. For $x>0$ we have
-$x^{r}=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}t,\qquad\operatorname{Re}(r)>-1.$
+$\int_{-\infty}^{+\infty}\frac{e^{-itx}}{(s+it)^{r+1}}\mathrm{d}t=0,\qquad\text{for}\quad x>0,\ s>0,\ \operatorname{Re}(r)>-1.$ (1)
+For $x=0$ the identity holds for $\operatorname{Re}(r)>0$.
+[^4] Obviously, for $x<0$ the identity is $\int_{-\infty}^{+\infty}\frac{e^{itx}}{(s+it)^{r+1}}\mathrm{d}t=0$ for $s>0$ and $\operatorname{Re}(r)>-1$.
+The identity in (1) is obtained by a contour argument that is valid for $\operatorname{Re}(r)>-1$. Note that $f\in L^{1}(\mathbb{R})$ if and only if $\operatorname{Re}(r)>0$, because
+$\left|f(t)\right|=\left|\frac{e^{-itx}}{(s+it)^{r+1}}\right|=\frac{1}{|s+it|^{\operatorname{Re}(r)+1}}=\frac{1}{(s^{2}+t^{2})^{(\operatorname{Re}(r)+1)/2}}.$
+For $1<\operatorname{Re}(r)\leq 0$ the integral is not absolutely convergent; in this range we interpret the formula (1) as the oscillatory improper integral,
+$\lim_{T\to\infty}\int_{-T}^{T}\frac{e^{-itx}}{(s+it)^{r+1}}\mathrm{d}t,$
+which exists and equals 0 for $x>0$ by closing the contour in the lower half-plane ($e^{-itx}$ decays there) and observing that $s+it\neq 0$ for $t\in\mathbb{R}$ when $s>0$.
+For the special case $x=0$ and $r=0$, it is easy to establish the well-known identity
+$\int_{-\infty}^{+\infty}\frac{1}{s+it}\mathrm{d}t=\pi\operatorname{sgn}(s),\qquad\text{for all }s\in\mathbb{R},$ (2)
+where $\operatorname{sgn}(s)\equiv\frac{s}{|s|}1_{\{s\neq 0\}}\in\{-1,0,1\}$.
+In this setting where $s>0$, we simply have $\int_{-\infty}^{+\infty}\frac{1}{s+it}\mathrm{d}t=\pi$, which we will use in the first formula of following Lemma.
+Lemma 2.
+Let $x\in\mathbb{R}$ and $z=s+it\in\mathbb{C}$, where $s>0$ is an arbitrary constant. Then
+$|x|^{r}=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}+e^{-zx}}{z^{r+1}}\mathrm{d}t,\qquad\text{for}\quad\operatorname{Re}(r)>-1,\quad x\neq 0,$
+and this identity also holds for $x=0$ if $\operatorname{Re}(r)>0$ or r=0r=0.555When $x^{r}$ is to be evaluated for $x=r=0$, we follow the standard convention $0^{0}=1$.
+For $x>0$ we have
+$x^{r}=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}t,\qquad\text{for}\quad\operatorname{Re}(r)>-1,\quad x>0,$
+and this identity also holds for $x=0$ if $\operatorname{Re}(r)>0$.
 Similarly,
-$x^{k}$ $=$ $\frac{k!}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}+\left(-1\right)^{k}e^{-zx}}{z^{k+1}}\mathrm{d}t,\qquad k\in\mathbb{N}_{0},$
-for any $x\neq 0$, and this identity also holds for $x=0$ if $k\in\mathbb{N}$.
-The case, $x=0$, arises as a special case in Lemma 1, requiring results to include a qualification for distributions with positive mass at zero.
-Our first result is the following expression for fractional absolute (possibly complex) moments.
+$x^{k}$ $=$ $\frac{k!}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}+\left(-1\right)^{k}e^{-zx}}{z^{k+1}}\mathrm{d}t,\qquad k\in\mathbb{N}_{0},\quad x\in\mathbb{R}.$
+Here $x=0$, arises as a special case in the first two formulas of Lemma 2, and this requiring results to include a qualification for distributions with positive mass at zero.
+Several results will now follows by taking expectation to both sides of the identities in Lemma 2, and justify interchanging expectation and integral.
+For $\operatorname{Re}(r)>0$ this is straight forward by means of Fubini/Tonelli theorems.
+However, as detailed above, the integrand is not absolute integrable for $\operatorname{Re}(r)\leq 0$, and interchanging expectation and integral requires a different line of arguments, and the improper integrals in the form $\int_{-\infty}^{+\infty}[\cdot]\mathrm{d}t$ is to be interpreted as $\lim_{T\rightarrow\infty}\int_{-T}^{+T}[\cdot]\mathrm{d}t$ for $\operatorname{Re}(r)\in(-1,0]$.
+Our first main result concerns fractional absolute moments.
 Theorem 1 (Absolute moments).
-Suppose $X$ satisfies Assumption 1, $\xi\in\mathbb{R}$, and $z=s+it$ with $s\in\mathcal{N}_{+}$. Then
-$\mathbb{E}|X-\xi|^{r}$ $=$ $\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-\xi z}M_{X}(z)+e^{\xi z}M_{X}(-z)}{z^{r+1}}\mathrm{d}t,$ (1)
-for $\operatorname{Re}(r)\geq 0$, and if $\Pr(X=\xi)=0$, then the identity holds for all $\operatorname{Re}(r)>-1$.
-If $r$ is real, then (1) is equivalent to
-$\mathbb{E}|X-\xi|^{r}=\frac{\Gamma(r+1)}{\pi}\int_{0}^{\infty}\operatorname{Re}\left[\frac{e^{-\xi z}M_{X}(z)+e^{\xi z}M_{X}(-z)}{z^{r+1}}\right]\mathrm{d}t.$ (2)
-The result obviously covers regular and central absolute moments by setting $\xi=0$ for regular moments and $\xi=\mathbb{E}X$ for central moments.
-Note that the expression remains invariant to the specific value of $s$, so long as $s\in\mathcal{N}_{+}$.
-In our experience, the choice of $s$ has no practical impact on moment evaluation.
+Suppose that $\mathbb{E}[e^{\pm sX}]<\infty$ for some $s>0$. For $\operatorname{Re}(r)\in(-1,0)$, assume additionally that $\mathbb{E}|X-\xi|^{r}<\infty$. If $\Pr(X=\xi)=0$ then
+$\mathbb{E}|X-\xi|^{r}=\frac{\Gamma(r+1)}{\pi}\int_{0}^{\infty}\operatorname{Re}\left[\frac{e^{-\xi z}M_{X}(z)+e^{\xi z}M_{X}(-z)}{z^{r+1}}\right]\mathrm{d}t,\quad\text{for }r>-1,$ (3)
+and for complex moments, $r\in\mathbb{C}$, we have
+$\mathbb{E}|X-\xi|^{r}=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-\xi z}M_{X}(z)+e^{\xi z}M_{X}(-z)}{z^{r+1}}\mathrm{d}t,\quad\operatorname{Re}(r)>-1.$ (4)
+Moreover, if $\Pr(X=\xi)>0$, the identities hold for $r\geq 0$ and $\operatorname{Re}(r)>0$ respectively.
+The finite moment requirement, $\mathbb{E}|X-\xi|^{r}<\infty$, is implied by $\mathbb{E}[e^{\pm sX}]<\infty$ for $\operatorname{Re}(r)>0$, but is not guaranteed for $\operatorname{Re}(r)<0$.
+The results in Theorem 1 include regular absolute moments by setting $\xi=0$ and central moments by setting $\xi=\mathbb{E}X$.
 Although analytical integration may often be impractical, (4) offers a novel method to evaluate moments numerically.
 For real moments, $r\in\mathbb{R}$, it is our experience that it is computationally advantageous to use (3) rather than (4).
+Remark.
+$\mathbb{E}[e^{\pm sX}]<\infty$ for some $s>0$ is equivalent to assuming the MGF is well defined in a neighborhood of zero.
+In next Theorem we only assume $\mathbb{E}[e^{sX}]<\infty$ for some $s>0$, which does not guarantee the MGF is well-defined in a neighborhood about zero.
 Interestingly, Theorem 1 can be used to obtain new identities, by equating (4) with an existing expression for absolute moments.
-For instance, the absolute moment for standard normal random variable, $X\sim N(0,1)$, is $\mathbb{E}|X|^{r}=\frac{\Gamma(\frac{r+1}{2})}{\sqrt{\pi}}2^{r/2}$, for $r>-1$, and equating this with (1) yields the following expression for the reciprocal gamma function
+For instance, the absolute moment of a standard normal random variable, $X\sim N(0,1)$, is $\mathbb{E}|X|^{r}=\frac{\Gamma(\frac{r+1}{2})}{\sqrt{\pi}}2^{r/2}$, for r>1r>-1, and equating this with (4) yields the following expression for the reciprocal gamma function
 $\frac{1}{\Gamma(\frac{r}{2}+1)}=\frac{2^{r/2}}{\pi}\int_{-\infty}^{+\infty}\frac{e^{z^{2}/2}}{z^{r+1}}\mathrm{d}t,\qquad r>-1,$ (5)
-for $z=s+it$, where $s>0$ is an arbitrary positive constant. While we can only infer that (3) holds for real $r>-1$, the identity actually holds for any $r\in\mathbb{C}$, see Hansen and Tong, (2024).
-For non-negative random variables we have the following simpler expression for fractional moments.
-Theorem 2 (Moments of non-negative random variables).
-Let $X$ satisfies Assumption 1 and $\Pr(X\geq\xi)=1$ for some $\xi\in\mathbb{R}$. For any $r\in\mathbb{C}$ with $\operatorname{Re}(r)\geq 0$, we have,
-$\mathbb{E}[(X-\xi)^{r}]=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-\xi z}M_{X}(z)}{z^{r+1}}\mathrm{d}t,$ (4)
-where $z=s+it$ with $s\in\mathcal{N}_{+}$. If $X$ is strictly greater than $\xi$, i.e. $\Pr(X>\xi)=1$, then the identity holds for $\operatorname{Re}(r)>-1$.
-If $r$ is real, then (4) is equivalent to
-$\mathbb{E}[(X-\xi)^{r}]=\frac{\Gamma(r+1)}{\pi}\int_{0}^{\infty}{\rm Re}\left[\frac{e^{-\xi z}M_{X}(z)}{z^{r+1}}\right]\mathrm{d}t.$ (5)
+for $z=s+it$, where $s>0$ is an arbitrary positive constant.
+From Theorem 1 we can only infer that (5) holds for real r>1r>-1, but the identity actually holds for any $r\in\mathbb{C}$, such that we obtain a new holomorphic integral representation of Gamma function and its reciprocal that does not rely on analytic continuation, see Hansen and Tong, (2025).
+Below we will derive an expression for positive-part moments, where we use the convention $x_{+}^{r}=x^{r}1_{\{x>0\}}$.
+This enables us to accommodate negative powers, because $0_{+}^{r}=0$ for all rr.
+[^6] Note that with this convention we have ($r=0$) we have $0=0_{+}^{0}$, unlike $0^{0}=1$).
+Importantly, for $\operatorname{Re}(r)\in(-1,0]$, the integral $\int_{-\infty}^{+\infty}\frac{e^{-\xi z}M_{X}(z)}{z^{r+1}}\mathrm{d}t$, is to be interpreted as the symmetric Cauchy principal value, $\lim_{T\rightarrow\infty}\int_{-T}^{+T}\frac{e^{-\xi z}M_{X}(z)}{z^{r+1}}\mathrm{d}t$.
+Theorem 2 (Positive-part moments).
+Suppose that $\mathbb{E}[e^{sX}]<\infty$ for some $s>0$. For $\operatorname{Re}(r)\in(-1,0)$, assume additionally that $\mathbb{E}|X-\xi|^{r}<\infty$. Then
+$\mathbb{E}[(X-\xi)_{+}^{r}]=\frac{\Gamma(r+1)}{\pi}\int_{0}^{\infty}{\rm Re}\left[\frac{e^{-\xi z}M_{X}(z)}{z^{r+1}}\right]\mathrm{d}t,\quad\text{for }r>0,$ (6)
+and the the identity also holds for $r\in(-1,0]$ if $\Pr(X=\xi)=0$.
+For complex moments, $r\in\mathbb{C}$, we have
+$\mathbb{E}[(X-\xi)_{+}^{r}]=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-\xi z}M_{X}(z)}{z^{r+1}}\mathrm{d}t,\quad\text{for }\operatorname{Re}(r)>-1.$ (7)
+Note that $\mathbb{E}[e^{sX}]<\infty$ for some $s>0$ ensures that $\mathbb{E}(X-\xi)_{+}^{r}<\infty$ for ${\rm Re}\left(r\right)>0$ but not for $\operatorname{Re}(r)\in\left(-1,0\right)$.
+Therefore, for $\operatorname{Re}(r)\in\left(-1,0\right)$ where we rely on the dominated convergence theorem, the condition $\mathbb{E}|X-\xi|^{r}<\infty$ is required for interchanging expectation and integral.
+Remark 1.
+Here we assume $\mathbb{E}[e^{sX}]<\infty$ for some $s>0$, which does not guarantee the existence of the MGF in an open neighborhood of zero.
+So, $\mathbb{E}|X-\xi|^{r}<\infty$ is not guaranteed by $\mathbb{E}[e^{sX}]<\infty$, as illustrated with the example below.
+Note that $\mathbb{E}[e^{sX}]<\infty$ for some $s>0$ does ensure $\mathbb{E}(X-\xi)_{+}^{r}<\infty$ for $\operatorname{Re}(r)>0$.
+Example 1 (One-sided MGF and divergent moments.).
+Let $Y\sim\mathrm{Exp}(1)$, $Z\sim\mathrm{Pareto}(\alpha)$ with $\mathbb{P}(Z>z)=z^{-\alpha}$ for $z\geq 1$, and choose $0<\alpha<1$. Define
+$X=\begin{cases}+Y,&\text{with prob. }1/2,\\ -Z,&\text{with prob. }1/2,\end{cases}$
+with $Y$, $Z$ independent.
+Then, for $0<s<1$, $\mathbb{E}[e^{sX}]=\tfrac{1}{2}\mathbb{E}[e^{sY}]+\tfrac{1}{2}\mathbb{E}[e^{-sZ}]=\tfrac{1}{2}(1-s)^{-1}+\tfrac{1}{2}\mathbb{E}[e^{-sZ}]<\infty$.
+However, for $\tilde{s}<0$ we have $\mathbb{E}[e^{\tilde{s}X}]\geq\tfrac{1}{2}\mathbb{E}[e^{|\tilde{s}|Z}]=\tfrac{1}{2}\int_{1}^{\infty}e^{|\tilde{s}|z}\alpha z^{-\alpha-1}\mathrm{d}z=\infty$.
+Hence the MGF is not finite on any open neighborhood of 0.
+For $r\in(-1,\alpha)$ we have $\mathbb{E}|X|^{r}=\tfrac{1}{2}\Gamma(r+1)+\tfrac{1}{2}\frac{\alpha}{\alpha-r}$, while for $r\geq\alpha$, $\mathbb{E}[|X|^{r}]\geq\tfrac{1}{2}\mathbb{E}[Z^{r}]=\infty$.
+We can nevertheless apply Theorem 2 for r>1r>-1 because $\mathbb{E}(X)_{+}^{r}=\tfrac{1}{2}\Gamma(r+1)$.
+The expression in Theorem 2 is identical to one in Pinelis, (2011, theorem 1), albeit we extend the range of moments to include some negative moments, $r\in(-1,0)$, and complex moments with $\operatorname{Re}(r)>-1$.
+The special case $r=1$ has received considerable attention because $\mathbb{E}[(X-\xi)_{+}]$ plays a central role in option pricing, with $\xi$ representing the strike price.
+For this special case, $r=1$, (7) coincides with Kim et al., (2010, lemma 7.1) and (6) coincides with Huang and Oosterlee, (2011, eq.
+3.1).
+[^7] They derived the result under slightly stronger assumptions.
+In Kim et al., (2010) it is assumed that the distribution is infinitely divisible and continuous, while Huang and Oosterlee, (2011, eq.
+3.1) assume continuity of the distribution.
+Neither of these assumptions is required here.
+Note that the expression remains invariant to the value of $s>0$, and numerical integration is, in our experience, insensitive to the choice of $s$, so long as $\mathbb{E}[e^{sX}]<\infty$.
+For the positive-part moments, we do not require $\mathbb{E}[e^{uX}]<\infty$ for any $u<0$.
+For non-negative random variables the following is a simple implication of Theorem 2 and Lemma 2.
+Corollary 1 (Moments of nonnegative random variables).
+Suppose that $\mathbb{E}[e^{sX}]<\infty$ for some $s>0$ and for $\operatorname{Re}(r)\in(-1,0)$, assume additionally that $\mathbb{E}|X-\xi|^{r}<\infty$. (Strictly positive). If $\Pr(X>\xi)=1$, then
+$\mathbb{E}[(X-\xi)^{r}]=\frac{\Gamma(r+1)}{\pi}\int_{0}^{\infty}{\rm Re}\left[\frac{e^{-\xi z}M_{X}(z)}{z^{r+1}}\right]\mathrm{d}t\quad\text{for }r>-1.$ (8)
+(Nonnegative). If $\Pr(X\geq\xi)=1$, then the identity holds for $r>0$, and for complex moments, $r\in\mathbb{C}$, we have the identity
+$\mathbb{E}[(X-\xi)^{r}]=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-\xi z}M_{X}(z)}{z^{r+1}}\mathrm{d}t,\quad\text{for }\operatorname{Re}(r)>-1.$ (9)
+The condition $\mathbb{E}[(X-\xi)^{r}]<\infty$ is redundant for $\operatorname{Re}(r)>0$, because it is implied by $\mathbb{E}[e^{sX}]<\infty$, but for negative moments this additional conditions is needed.
 As an illustration, we apply (9) to an exponentially distributed random variable.
@@
 Example 2.
-For an exponentially distributed random variable, with parameter $\lambda>0$, $X\sim\operatorname{Exp}(\lambda)$, we have $M_{X}(z)=\frac{\lambda}{\lambda-z}$, with $\mathcal{N}=(-\infty,\lambda)$.
+For an exponentially distributed random variable, with parameter $\lambda>0$, $X\sim\operatorname{Exp}(\lambda)$, we have $M_{X}(z)=\frac{\lambda}{\lambda-z}$, with region of convergence $\operatorname{Re}(z)<\lambda$.
 So, the integral in (8) is $\int_{0}^{\infty}{\rm Re}\left[\frac{\lambda}{\lambda-z}\frac{1}{z^{r+1}}\right]\mathrm{d}t$.
@@
 $=\lambda^{-k}\int_{0}^{\infty}{\rm Re}\left[\tfrac{\lambda-z^{\ast}}{(\lambda-z^{\ast})(\lambda-z)}+\tfrac{z^{\ast}}{z^{\ast}z}\right]\mathrm{d}t,$
-because $\int_{0}^{\infty}{\rm Re}\left[z^{-j}\right]\mathrm{d}t=0$ for $j>1$. Substituting $z=s+it$ in the last expression gives us
+as $\int_{0}^{\infty}{\rm Re}\left[z^{-j}\right]\mathrm{d}t=0$ for $j>1$. Substituting $z=s+it$ in the last expression gives us
 $\lambda^{-k}\int_{0}^{\infty}\operatorname{Re}\left[\tfrac{\lambda-s+it}{(\lambda-s)^{2}-(it)^{2}}+\tfrac{s-it}{s^{2}-(it)^{2}}\right]\mathrm{d}t$ $=$ $\lambda^{-k}\int_{0}^{\infty}\left[\tfrac{\lambda-s}{(\lambda-s)^{2}+t^{2}}+\tfrac{s}{s^{2}+t^{2}}\right]\mathrm{d}t$
 $=$ $\lambda^{-k}\left[\left.\arctan(\tfrac{t}{\lambda-s})+\arctan(\tfrac{t}{s})\right|_{0}^{\infty}\right]=\pi\lambda^{-k}.$
-Thus that $\mathbb{E}[X^{k}]=\lambda^{-k}k!$ as expected.
-Deriving non-integer moments, $r\notin\mathbb{N}$, can be derived using contour integrals and the Cauchy Integral Theorem.
-These are the techniques used extensively to prove the main results in this paper.
+Thus that $\mathbb{E}[X^{k}]=\lambda^{-k}k!$ as expected. Non-integer moments, $r\notin\mathbb{N}$, can be derived using contour integrals and the Cauchy Integral Theorem.
 For random variables whose support may include negative numbers, we have the following result for integer moments.
 Theorem 3.
-Let $X$ satisfy Assumption 1 and let $\xi\in\mathbb{R}$. For $k\in\mathbb{N}$ we have
-$\mathbb{E}[(X-\xi)^{k}]$ $=$ $\frac{k!}{\pi}\int_{0}^{\infty}\operatorname{Re}\left[\frac{e^{-\xi z}M_{X}(z)+(-1)^{k}e^{\xi z}M_{X}(-z)}{z^{k+1}}\right]\mathrm{d}t,$ (6)
-where $z=s+it$ with $s\in\mathcal{N}_{+}$.
-Because $k$ is real, we have expressed the integral on the form, $\frac{k!}{\pi}\int_{0}^{\infty}\operatorname{Re}[\cdot]\mathrm{d}t$, which tends to be simpler to evaluate numerically than the equivalent expression using the form, $\frac{k!}{2\pi}\int_{-\infty}^{+\infty}[\cdot]\mathrm{d}t$.
+Suppose $\mathbb{E}[e^{\pm sX}]<\infty$ for some $s>0$. Then
+$\mathbb{E}[(X-\xi)^{k}]$ $=$ $\frac{k!}{\pi}\int_{0}^{\infty}\operatorname{Re}\left[\frac{e^{-\xi z}M_{X}(z)+(-1)^{k}e^{\xi z}M_{X}(-z)}{z^{k+1}}\right]\mathrm{d}t,\quad\text{for }k\in\mathbb{N}_{0}.$ (10)
+Because $k$ is real, we have expressed the integral in the form, $\frac{k!}{\pi}\int_{0}^{\infty}\operatorname{Re}[\cdot]\mathrm{d}t$, which tends to be simpler to evaluate numerically than the equivalent expression using the form, $\frac{k!}{2\pi}\int_{-\infty}^{+\infty}[\cdot]\mathrm{d}t$.
 Remark.
 In some situations, Theorems 1, 2, and 3, provide different expressions for the same moments.
-For instance, (4) and (1) must agree for a non-negative random variable, because $\mathbb{E}|X|^{r}=\mathbb{E}[X^{r}]$ in this case, and for integer moments, these expressions must also agree with (6).
+For instance, (4) and (9) must agree for a non-negative random variable, because $\mathbb{E}|X|^{r}=\mathbb{E}[X^{r}]$, and for integer moments, these expressions must also agree with (10).
 This is indeed the case, because the additional term in (4) and the additional term in (10) are both zero for non-negative variables.
-This is an implication of Lemma A.2, which establishes that $\int_{-\infty}^{\infty}\frac{e^{-\left(s+it\right)x}}{\left(s+it\right)^{r+1}}\mathrm{d}t=0$ for all $x>0$ and all $s>0$.
-We can combine Theorems 1 and 3 to compute odd moments over tail events.
-Corollary 1 (Tail Moments).
-Let $X$ satisfy Assumption 1, $\xi\in\mathbb{R}$, and let $k$ be odd number. Then
-$\mathbb{E}[(X-\xi)^{k}1_{\{X>\xi\}}]$ $=$ $+\frac{k!}{\pi}\int_{0}^{\infty}{\rm Re}\left[\frac{e^{-\xi z}M_{X}(z)}{z^{k+1}}\right]\mathrm{d}t,$
-$\mathbb{E}[(X-\xi)^{k}1_{\{X<\xi\}}]$ $=$ $-\frac{k!}{\pi}\int_{0}^{\infty}{\rm Re}\left[\frac{e^{-\xi z}M_{X}(z)}{z^{k+1}}\right]\mathrm{d}t,$
-where $z=s+it$ for any $s\in\mathcal{N}_{+}$.
-We can also compute the cdf, quantiles, and expected shortfall using the new moment expressions. We can obtain the cdf for a continuous random variable, $X$, from the MGF by Gil-Pelaez inversion,
-$F_{X}(x)=\frac{1}{2}-\frac{1}{\pi}\int_{0}^{\infty}{\rm Re}\left[\frac{M_{X}(it)e^{-itx}}{it}\right]\mathrm{d}t,$
-where we used that $M_{X}(it)$ is the characteristic function for $X$.
-Now the $\alpha$-quantile is given by $\xi_{\alpha}\equiv F_{X}^{-1}(\alpha)$, and we can use Corollary 1 (with $k=1$) to compute,
-$\operatorname{ES}_{\alpha}(X)=-\frac{1}{\alpha}\mathbb{E}[X1_{\{X<\xi_{\alpha}\}}]=-\frac{1}{\alpha}\mathbb{E}[(X-\xi_{\alpha})1_{\{X<\xi_{\alpha}\}}]-\xi_{\alpha},$
-which is known as the expected shortfall.
+This is an implication of Lemma 1, which establishes that $\int_{-\infty}^{\infty}\frac{e^{-\left(s+it\right)x}}{\left(s+it\right)^{r+1}}\mathrm{d}t=0$ for all $x>0$ and all $s>0$.
+Negative tail moments, which is part of the definition of expected shortfall (see below), have the following expression, where we can combine Theorems 2 and 3 and use the notation $x_{-}=\min(x,0)$.
+Corollary 2 (Negative tail moments).
+Suppose $\mathbb{E}[e^{-sX}]<\infty$ for some $s>0$ and $\mathbb{E}|X-\xi|^{k}<\infty$. Then
+$\mathbb{E}[(X-\xi)_{-}^{k}]=-\frac{k!}{\pi}\int_{0}^{\infty}{\rm Re}\left[\frac{e^{\xi z}M_{X}(-z)}{z^{k+1}}\right]\mathrm{d}t,\quad\text{for }k\text{ odd}.$
+We can also compute the cumulative distribution function (cdf), quantiles, and expected shortfall using the new moment expressions.
+For a continuous random variable, $X$, its cdf is explicitly given by, $F_{X}(x)=\frac{1}{2}-\frac{1}{\pi}\int_{0}^{\infty}{\rm Re}\left[\frac{M_{X}(it)e^{-itx}}{it}\right]\mathrm{d}t,$ which is the classical Gil-Pelaez inversion formula that provides a direct way to recover probability distributions from their transforms.
+Here we have used that $M_{X}(it)$ is the characteristic function for $X$.
+If unique, the $\alpha$-quantile is now given by $\xi_{\alpha}\equiv F_{X}^{-1}(\alpha)$, and we can use Corollary 2 with $k=1$ to compute expected shortfall,
+$\operatorname{ES}_{\alpha}(X)=-\frac{1}{\alpha}\mathbb{E}[X1_{\{X<\xi_{\alpha}\}}]=-\frac{1}{\alpha}\mathbb{E}[(X-\xi_{\alpha})1_{\{X<\xi_{\alpha}\}}]-\xi_{\alpha}.$
 2.1 Example: Moments of NIG Distribution
 To illustrate the new method, we use Theorem 1 to compute fractional moments of the normal-inverse Gaussian (NIG) distribution.
-The CMGF method is significantly faster and more accurate than both simulation-based methods and direct integration using the NIG density.
-The NIG distribution was introduced in Barndorff-Nielsen, (1978) and has four parameters, $\lambda$ (location), $\delta$ (scale), $\alpha$ (tail heaviness), and $\beta$ (asymmetry). It has density
-$f(x)=\tfrac{\alpha\delta}{\pi\sqrt{\delta^{2}+(x-\lambda)^{2}}}K_{1}\left(\alpha\sqrt{\delta^{2}+(x-\lambda)^{2}}\right)e^{\delta\gamma+\beta(x-\lambda)},\qquad x\in\mathbb{R}$
-where $\gamma=\sqrt{\alpha^{2}-\beta^{2}}$ and $K_{1}(\cdot)$ is the modified Bessel function of the second kind, and the MGF is given by $M_{X}(z)=\exp\left(\lambda z+\delta\left[\gamma-\sqrt{\alpha^{2}-(\beta-z)^{2}}\right]\right),$ for all $z\in\mathbb{C}$.
-We consider two standardized NIG distribution (zero mean and unit variance), which can be characterized by the two parameters, $(\xi,\chi)$, for $10\leq|\chi|<\xi<$ 1, where $\lambda=-\tfrac{\chi}{\xi}\zeta$, $\delta=\xi^{2}\sqrt{\xi^{2}-\chi^{2}}\zeta$, $\alpha=\xi\zeta$, and $\beta=\chi\zeta$, with $\zeta=\sqrt{1-\xi^{2}}/(\xi^{2}-\chi^{2})$, see Barndorff-Nielsen et al., (1985).
+The NIG distribution was introduced in Barndorff-Nielsen, (1978) and has four parameters, $\mu$ (location), $\delta$ (scale), $\alpha$ (tail heaviness), and $\beta$ (asymmetry).
+It has density
+$f(x)=\tfrac{\alpha\delta}{\pi\sqrt{\delta^{2}+(x-\mu)^{2}}}K_{1}\left(\alpha\sqrt{\delta^{2}+(x-\mu)^{2}}\right)e^{\delta\gamma+\beta(x-\mu)},\qquad x\in\mathbb{R}$
+where $\gamma=\sqrt{\alpha^{2}-\beta^{2}}$ and $K_{1}(\cdot)$ is the modified Bessel function of the second kind, and the MGF is given by $M_{X}(z)=\exp\left(\mu z+\delta\left[\gamma-\sqrt{\alpha^{2}-(\beta-z)^{2}}\right]\right),$ for $z=s+it$ with $\beta-\alpha<s<\beta+\alpha$.
+Existing expressions for fractional moments of NIG distributions are rather complicated.
+Let $X\sim\mathrm{NIG}(\alpha,\beta,\mu,\delta)$ and $Y\sim\mathrm{NIG}(\alpha,0,\mu,\delta)$, then Barndorff-Nielsen and Stelzer, (2005) showed that
+$\mathbb{E}|X|^{r}=e^{\delta(\gamma-\alpha)}\sum_{k=0}^{\infty}\frac{\beta^{k}}{k!}\mathbb{E}\left(|Y|^{r}(Y-\mu)^{k}\right),\qquad r>0,$
+which expresses $\mathbb{E}|X|^{r}$ as an infinite sum of moments involving the symmetric NIG, $Y$.
+For $\mu$-centered moments Barndorff-Nielsen and Stelzer, (2005, corollary 3) obtained the following explicit formula,
+$\mathbb{E}(|X-\mu|^{r})=\frac{\alpha}{\pi}\left(\frac{2\delta}{\alpha}\right)^{\tfrac{r+1}{2}}e^{\delta\gamma}\sum_{k=0}^{\infty}\frac{2^{k}\Gamma(k+\tfrac{r+1}{2})}{(2k)!}\left(\frac{\delta\beta^{2}}{\alpha}\right)^{k}K_{k+\tfrac{r-1}{2}}(\delta\alpha),\quad r>0.$ (11)
+The corresponding CMGF integral is
+$\mathbb{E}|X-\mu|^{r}=\frac{\Gamma(r+1)}{2\pi}e^{\delta\gamma}\int_{-\infty}^{+\infty}\frac{e^{-\delta\sqrt{\alpha^{2}-(\beta-z)^{2}}}+e^{-\delta\sqrt{\alpha^{2}-(\beta+z)^{2}}}}{z^{r+1}}\mathrm{d}t,\qquad r>-1,$
+and Theorem 1 yields similar expressions for any other recentering ($\xi\neq\mu$), which is not possible with (11).
+Evaluating the CMGF integral is also substantially faster and more accurate than both simulation-based methods and direct integration using the NIG density.
+To illustrate this, we consider two standardized NIG distribution (zero mean and unit variance), which can be characterized by the two parameters, $(\xi,\chi)$, for $0\leq|\chi|<\xi<1$, where $\mu=-\tfrac{\chi}{\xi}\zeta$, $\delta=\xi^{2}\sqrt{\xi^{2}-\chi^{2}}\zeta$, $\alpha=\xi\zeta$, and $\beta=\chi\zeta$, with $\zeta=\sqrt{1-\xi^{2}}/(\xi^{2}-\chi^{2})$, see Barndorff-Nielsen et al., (1985).
 Specifically we consider $(\xi,\chi)=(1/2,-1/3)$ and $(\xi,\chi)=(1/8,-1/16)$, the corresponding densities are shown in the left panel of Figure 1.
-The corresponding MGFs, $t\mapsto\operatorname{MGF}(s+it)$ are shown in the right panel for $s=1$ – solid lines for the real part and dashed lines for the complex part.
-[…]
-We compute the absolute moments for $-0.85\leq r\leq 4.2$ for both distributions using Theorem 1 and the moments are shown in the upper panel of Figure 2.
+The suitable range for $s$ is $0<s<\alpha+\beta$, where the upper limit for these distributions is $3\sqrt{3}/5\approx 1.04$ and $2\sqrt{63}/3\approx 5.29$, respectively.
+So, we can set $s=1$ and the mapping, and $t\mapsto\operatorname{MGF}(1+it)$ is shown for the two distributions in the right panel – solid lines for the real part and dashed lines for the imaginary part.
+Figure 1: Densities of two standardized NIG distributions (left panel) and the $t\mapstoMGF$($1+it)$ (right panel).
+The two distributions are for $(\xi,\chi)=(1/2,-1/3)$ (blue lines) and $(\xi,\chi)=(1/8,-1/16)$ (red lines).
+We compute the absolute moments for $0.85\leq r\leq 4.2$ for the two distributions using Theorem 1 and the moments are shown in the upper panel of Figure 2.
 The solid black dot is the second (absolute) moment, which is known to be one for these standardized distributions, and the solid blue and red dots represent the fourth (absolute) moment, which are known to equal $3(1+4\chi^{2})/(1-\xi^{2})$.
@@
 Figure 2: Absolute moments of the standardized NIG distribution with $(\xi,\chi)=(1/2,-1/3)$ (blue lines) and $(\xi,\chi)=(1/8,-1/16)$ (red lines).
-These are shown in the left panels, where dots are represent the known moments for $r=2$ and $r=4$.
-The middle panel is a snippet of the left panel which can be compared to simulation-based moments for 15 values of $r$.
-For each $r$ we computed $\mathbb{E}|X|^{r}$ 100 times as represented by the x-crosses.
-Each of these simulation-based moments are based on $N=$1,000,000 draws of NIG distribution.
+These are shown in the left panels, where dots represent the known moments for $r=2$ and $r=4$.
+The middle panel is a snippet of the left panel where we have added x-crosses that represents 100 simulation-based estimates of $\mathbb{E}|X|^{r}$, for $r\in\{3.82,3.84,\ldots,4.18\}$.
+Each estimate is based on N=N=1,000,000 random draws of the NIG distribution.
 The right panel shows how many decimal places are accurate with a one standard deviation simulation error.
-Additionally, we compute the absolute moments using $N=1,000,000$ independent draws from the two NIG distributions.
-The simulation based moment is given by $\hat{\mu}_{N}^{r}=\frac{1}{N}\sum_{j=1}^{N}|X_{i}|^{r}$ and we estimate the $\sigma^{2}(r)=\operatorname{var}(|X_{j}|^{r})$ with 100 million draws.
-The accuracy of a simulated moment can be computed from its standard error.
-Assuming an error equal to one standard deviation, then the number of accurate decimal places is given by $-\log_{10}(\sigma(r)/\sqrt{N})$.
-[^4] For instance, an error where the first three decimal places are zero, say $0.00099$, will map to $-\log_{10}(0.99\cdot 10^{-3})\simeq 3.004$.
+Additionally, we estimate the absolute moments using $N=1,000,000$ independent draws from the two NIG distributions.
+The simulation-based moment is given by $\hat{\mu}_{N}^{r}=\frac{1}{N}\sum_{j=1}^{N}|X_{i}|^{r}$ and we estimate the $\sigma^{2}(r)=\operatorname{var}(|X_{j}|^{r})$ with 100 million draws.
+The accuracy of a simulated moment is quantified by its standard error.
+Assuming an error equal to one standard deviation, then the number of accurate decimal places is given by $\log_{10}(\sigma(r)/\sqrt{N})$.
+[^8] For instance, a standard error equal to $0.00099$ and $N=10^{6}$ will map to $\log_{10}(0.99\cdot 10^{-3})\simeq 3.004$.
 We have plotted this number for $N=10^{6}$ in the lower left panel of Figure 2.
@@
 In the lower right panels we have shown the moments based on the CMGF method for $3.8\leq r\leq 4.2$ represented by the solid blue line.
-The $\times$-crosses are simulation-based moment estimates, each based on $N=1,000,000$ random draws.
+Each $\times$-cross is a simulation-based moment estimate based on $N=1,000,000$ random draws.
 Table 1: CMGF Computation Time of $\mathbb{E}|X|^{r}$ for $X\sim\operatorname{NIG}$
-$r$ CMGF Integrate with density Simulations $(N=10^{6})$
+$r$ CMGF $M_{X}^{(k)}$ Integrate with density Simulations $(N=10^{6})$
 -0.5 28.3 712.9 397,666
@@
 1.5 29.3 175.0 399,173
-2.0 20.4 108.3 298,635
+2.0 20.4 55.7 108.3 298,635
 2.5 24.5 149.3 399,182
@@
 3.5 24.5 114.7 402,552
-4.0 17.6 100.3 318,081
-Note: Computation time (in microseconds) for evaluating $\mathbb{E}|X|^{r}$ using three methods: the CMGF method of Theorem 1, by numerical integration of $\int_{-\infty}^{\infty}|x|^{r}f(x)\mathrm{d}x$, and by simulations, where we generate $X_{1},\ldots,X_{N}$ independent and identically NIG distributed, with $N=$1,000,000 and take the average of $|X_{i}|^{r}$.
+4.0 17.6 104.9 100.3 318,081
+Note: Computation time in microseconds (µs) for evaluating $\mathbb{E}|X|^{r}$ using four methods: the CMGF method of Theorem 1, the $k$-th derivative of the MGF (for $r=2$ and $r=4$), by numerical integration of $\int_{-\infty}^{\infty}|x|^{r}f(x)\mathrm{d}x$, and by simulations, where we generate $X_{1},\ldots,X_{N}$ independent and identically NIG distributed, with N=N=1,000,000 and take the average of $|X_{i}|^{r}$.
 Both CMGF and the density-based method use numerical integration with the same tolerance threshold.
 Computation times are evaluated with BenchmarkTools.jl for Julia, see Chen and Revels, (2016).
-Computation were done with Julia v1.11.0, see Bezanson et al., (2017), on a MacBook Pro M1 Max with 32 GM memory.
+Computations were done with Julia v1.11.0, see Bezanson et al., (2017), on a MacBook Pro M1 Max with 32 GM memory.
 The CMGF method is not only far more accurate in this example, it is also much faster than the simulation-based approach.
@@
 The CMGF method is between 5 and 25 times faster than the standard approach, and the CMGF method is more than 10,000 times faster than averaging (and generating) $N=10^{6}$ independent draws from the NIG distribution.
-We have also compared the methods using a standard normal distribution, for which the true moment is known to be, $\mathbb{E}|X|^{r}=\frac{\Gamma(\frac{r+1}{2})}{\sqrt{\pi}}2^{r/2}$, for all $r>-1$.
+We have also compared the methods using a standard normal distribution, for which the true moment is known to be, $\mathbb{E}|X|^{r}=\frac{\Gamma(\frac{r+1}{2})}{\sqrt{\pi}}2^{r/2}$, for all r>1r>-1.
 For the Gaussian case the CMGF is also thousands of times faster than simulation methods using 1 million draws and far more accurate.
@@
 $h_{t+1}$ $=\omega+\beta h_{t}+\alpha\left(z_{t}-\gamma\sqrt{h_{t}}\right)^{2},$
-where $r_{f}$ is the risk free rate; $\lambda$ is the equity risk premium; and $h_{t+1}={\rm var}\left(R_{t+1}|\mathcal{F}_{t}\right)$ is the conditional variance of daily log-returns $r_{t+1}=\log\left(S_{t+1}/S_{t}\right)$.
-Among the many variations of GARCH models, the HNG model stands out as having an analytical MGF for cumulative returns, $R_{T,H}=\sum_{t=T+1}^{T+H}r_{t}$.
+where $r_{f}$ is the risk-free rate; $\lambda$ is the equity risk premium; and $h_{t+1}={\rm var}\left(R_{t+1}|\mathcal{F}_{t}\right)$ is the conditional variance of daily log-returns $r_{t+1}=\log\left(S_{t+1}/S_{t}\right)$.
+Among the many variations of GARCH models, the HNG model stands out as having an analytical MGF for cumulative returns, $R_{T,H}=\sum_{t=T+1}^{T+H}r_{t}$, where $H$ is the horizon in trading days.
 The underlying reason is that its dynamic structure is carefully crafted for the purpose of yielding closed-form option pricing formulae, and option pricing formulae depend on the properties of cumulative returns, $R_{T,H}=\sum_{t=T+1}^{T+H}r_{t}$.
@@
 Proposition 1.
-Let $r_{t}$, $t=1,2,\ldots$, be given by (7) and define cumulative returns, $R_{T,H}=r_{T+1}+\cdots+r_{t+H}$. Then the conditional MFG has the affine form,
+Let $r_{t}$, $t=1,2,\ldots$, be given by (12) and define cumulative returns, $R_{T,H}=r_{T+1}+\cdots+r_{t+H}$. Then the conditional MGF has the affine form,
 $M_{R_{T,H}|T}(z)=\exp\left(A(H,z)+B(H,z)h_{T+1}\right),$ (13)
@@
 We will illustrate this with a simulation design that is based on real data.
-We estimate the HNG model using daily log returns for the S&P 500 index from January 1, 2000 to December 30, 2021.
+We estimate the HNG model using daily log-returns for the S&P 500 index from January 1, 2000 to December 30, 2021.
 The data were obtained from WRDS and the maximum likelihood estimates are presented in Table 2 along with their standard errors (in parentheses).
 Table 2: Heston-Nandi GARCH Estimation Results for Daily Log-returns of SPX
+$\lambda$ $\omega$ $\beta$ $\alpha$ $\gamma$ $\ell$
 […]
-1.9781 1.15 ×$\times 10^{-14}$ 0.7593 5.67×$\times 10^{-6}$ 185.5 17,963
-(0.3166) (0.85×$\times 10^{-14}$) (0.0190) (1.18×$\times 10^{-6}$) (23.5)
+[…]
 Note: Maximum likelihood estimates of the Heston-Nandi GARCH model based on daily S&P 500 returns with robust standard errors in parentheses. Sample period: January 1, 2000 to December 30, 2021.
@@
 The inverse of volatility corresponds to the negative moment, $r=-\frac{1}{2}$.
-This moment is interesting because it is used for Shape ratio forecasting and asset allocation, where the inverse of covariance matrix is employed.
+This moment is interesting because it is used for Sharpe ratio forecasting and asset allocation, where the inverse of covariance matrix is employed.
 We compute these four moments by Theorem 1.
@@
 Table 3: HARG Estimation Results for Daily Realized Variance
-[…]
+$\tilde{\beta}_{d}$ $\tilde{\beta}_{w}$ $\tilde{\beta}_{m}$ $\eta$ $\delta$ $\ell$
 0.4896 0.2789 0.0357 0.0053 0.9644 17,298
@@
 Table 4: ARP Model Estimation Results for Count data of CBOE VIX Jumps
-[…]
+$\omega$ $\beta$ $\alpha$ $\mathbb{E}(\lambda)$ $\ell$
 0.1548 0.7473 0.2043 3.2024 -9722
@@
 Appendix A Appendix of Proofs
-Some notation: We write +=$\mathbb{C}^{+}=\{z\in\mathbb{C}:\mathrm{Re}[z]>0\}$ to denote the complex numbers with positive real part.
-In the proof of Theorem 1, we will need the following result (the Laplace transformation of $x^{r}$).
 Lemma A.3.
-Let $r+1\in\mathbb{C}^{+}$ and $z\in\mathbb{C}^{+}$. Then
-$\int_{0}^{\infty}x^{r}e^{-zx}\mathrm{d}x$ $=\frac{\Gamma(r+1)}{z^{r+1}}.$ (A.1)
+For $x>0$ and $\operatorname{Re}(r)\in(-1,0)$
+$\int_{0}^{\infty}\frac{e^{\pm itx}}{t^{r+1}}\mathrm{d}t=x^{r}\Gamma(-r)e^{\mp\frac{i\pi r}{2}}.$
 Proof.
-Consider the case where $z$ is real and positive, i.e.
-$z=s>0$.
-From the definition of the gamma function and the substitution, $y\mapsto sx$, we have $\Gamma(r+1)=\int_{0}^{\infty}y^{r}e^{-y}\mathrm{d}y=\int_{0}^{\infty}s^{r}x^{r}e^{-sx}s\mathrm{d}x$, which proves the identity for any $z\in\mathbb{R}^{+}$.
-Both sides of (A.1) are holomorphic functions for $z\in\mathbb{C}^{+}$, and since they agree on $z\in\mathbb{R}^{+}$, it follows by the identity theorem that they agree for all $z\in\mathbb{C}^{+}$.
-[^9] We thank Christian Berg for suggesting this simplification of the proof.
-A different, direct proof for all $z\in\mathbb{C}^{+}$, is presented in the Supplementary Material.
-∎
-Lemma A.2.
-For $x<0$, $\operatorname{Re}(z)>0$, and $\operatorname{Re}(r)>-1$ we have
-$\int_{-\infty}^{+\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}t=0.$
-For $x=0$, the identity holds for $\operatorname{Re}(r)>0$.
-Proof.
-We first consider the case $x=0$. The result follows directly from,
-$\int_{-\infty}^{+\infty}\frac{e^{z\cdot 0}}{z^{r+1}}\mathrm{d}t=\int_{-\infty}^{+\infty}\frac{1}{(s+it)^{r+1}}\mathrm{d}t=\left.\frac{i}{r}\frac{1}{(s+it)^{r}}\right|_{-\infty}^{+\infty},$
-because the last term is zero for ${\rm Re}\left(r\right)>0$.
-Next, for $x<0$ we have
-$\int_{-\infty}^{+\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}t$ $=\int_{-\infty}^{+\infty}\frac{e^{\left(s+it\right)x}}{\left(s+it\right)^{r+1}}\mathrm{d}t$
-$=-\int_{+\infty}^{-\infty}\frac{e^{\left(s-i\tau\right)x}}{\left(s-i\tau\right)^{r+1}}\mathrm{d}\tau,\quad\tau=-t$
-$=\int_{-\infty}^{+\infty}\frac{e^{\left(s-i\tau\right)x}}{\left(s-i\tau\right)^{r+1}}\mathrm{d}\tau$
-$=e^{sx}\int_{-\infty}^{+\infty}\frac{e^{-i\tau x}}{\left(s-i\tau\right)^{r+1}}\mathrm{d}\tau$
-$=\lim_{R\rightarrow\infty}e^{sx}\int_{-R}^{+R}f(\tau)\mathrm{d}\tau,$
-where $f(\tau)=\frac{e^{-i\tau x}}{\left(s-i\tau\right)^{r+1}}$. The seemingly odd change of variables, $\tau=-t$, serves two important purposes that we explain below.
-We will use a closed contour integral in the complex plane with two part.
-The first part is the line segment along the real axis, from $-R$ to $R$, and the second part is the semicircular arc, $C_{R}=\{Re^{i\theta},\theta\in[0,\pi]\}$.
-The integral over the line segment, $\int_{-R}^{+R}f(\tau)\mathrm{d}\tau$, is the object of interest, whereas the semicircle defines an auxiliary integral that is used to close the contour, such that we can apply the Cauchy Integral Theorem.
-The first reason for the change of variable, $\tau=-t$, is to place the pole of $f(\zeta)$ outside the closed contour for all $R>0$ (the pole is at $\zeta=-is$), such that $0=\int_{-R}^{R}f(\zeta)\mathrm{d}\zeta+\int_{C_{R}}f(\zeta)\mathrm{d}\zeta$ by the Cauchy Integral Theorem.
+From Gradshteyn and Ryzhik, (2007, 3.761.4 and 3.761.9) we have
+$\int_{0}^{\infty}t^{\mu-1}\cos(tx)\mathrm{d}t$ $=$ $\frac{\Gamma(\mu)}{x^{\mu}}\cos\left(\frac{\pi\mu}{2}\right),\quad x>0,\operatorname{Re}(\mu)\in(0,1),$
+$\int_{0}^{\infty}t^{\mu-1}\sin(tx)\mathrm{d}t$ $=$ $\frac{\Gamma(\mu)}{x^{\mu}}\sin\left(\frac{\pi\mu}{2}\right),\quad x>0,\left|\operatorname{Re}(\mu)\right|\in(0,1).$
+Using $e^{\pm itx}=\cos(tx)\pm i\sin(tx)$, we have
+$\int_{0}^{\infty}t^{\mu-1}e^{\pm itx}\mathrm{d}t$ $=$ $\int_{0}^{\infty}t^{\mu-1}[\cos(tx)\pm i\sin(tx)]\mathrm{d}t$
+$=$ $\frac{\Gamma(\mu)}{x^{\mu}}\cos\left(\frac{\pi\mu}{2}\right)\pm i\frac{\Gamma(\mu)}{x^{\mu}}\sin\left(\frac{\pi\mu}{2}\right)=\frac{\Gamma(\mu)}{x^{\mu}}e^{\pm\frac{i\pi\mu}{2}}.$
+The result now follows by substituting $\mu=-r$, and noting that $\operatorname{Re}(\mu)\in(0,1)$ for $\operatorname{Re}(r)\in(-1,0)$. ∎
+Proof of Lemma 1. Observe that
+$\int_{-\infty}^{+\infty}\frac{e^{-itx}}{(s+it)^{\nu}}\mathrm{d}t=\int_{-\infty}^{+\infty}\frac{e^{i\tau x}}{(s-i\tau)^{\nu}}\mathrm{d}\tau=\lim_{R\rightarrow\infty}\int_{-R}^{+R}f(\tau)\mathrm{d}\tau,$
+where $f(\tau)=\frac{e^{i\tau x}}{(s-i\tau)^{\nu}}$. Here, we applied the change of variables, $\tau=-t$, for two purposes, which will be explained below.
+Let $x>0$ and consider a closed contour integral that consists of the line segment along the real axis, from R-R to $R$, and the semicircular arc, $C_{R}=\{Re^{i\theta},\theta\in[0,\pi]\}$.
+Our object of interest is the integral over the line segment, $\int_{-R}^{+R}f(\tau)\mathrm{d}\tau$, whereas the semicircle defines an auxiliary integral used to close the contour.
+The first reason for the change of variable, $\tau=-t$, is to ensure that the pole of $f(\zeta)$ (located at $\zeta=-is$) lies outside the closed contour for all $R>0$ because $s>0$.
+Thus, by the Cauchy Integral Theorem, $0=\int_{-R}^{R}f(\zeta)\mathrm{d}\zeta+\int_{C_{R}}f(\zeta)\mathrm{d}\zeta$.
 For the integral along the arc, we have the expression,
-$\int_{C_{R}}f(\zeta)\mathrm{d}\zeta=\int_{C_{R}}e^{-i\zeta x}g(\zeta)\mathrm{d}\zeta=\int_{C_{R}}e^{ia\zeta}g(\zeta)\mathrm{d}\zeta,$
-where $a=-x>0$ and $g(\zeta)=\frac{1}{(s-i\zeta)^{r+1}}$ is continuous and satisfies $\lim_{|\zeta|\rightarrow\infty}g(\zeta)=0$ for all $r+1\in\mathbb{C}^{+}$.
-The second reason for the change of variable is that it enables us to apply Jordan’s lemma, which requires $a>0$ in the expression above.
-By Jordan’s lemma we have that $\lim_{R\rightarrow\infty}\int_{C_{R}}e^{-ix\zeta}g(\zeta)\mathrm{d}\zeta=0$, which implies that
-$\int_{-\infty}^{+\infty}f(\tau)\mathrm{d}\tau=\lim_{R\rightarrow\infty}\int_{-R}^{+R}f(\tau)\mathrm{d}\tau=-\lim_{R\rightarrow\infty}\int_{C_{R}}f(\zeta)\mathrm{d}\zeta=0.$
-This completes the proof. ∎
-Proof of Lemma 1. Consider first the case $x\geq 0$. Since $\mathrm{Re}[z]=s>0$, we have the following identity from Lemma A.1,
-$\int_{0}^{\infty}x^{r}e^{-zx}\mathrm{d}x$ $=$ $\frac{\Gamma(r+1)}{z^{r+1}},\qquad\text{for all}\quad\mathrm{Re}[r]>-1.$
-This is the Laplace transform of $x^{r}$ and the inverse Laplace transform is
-$x^{r}=\frac{1}{2\pi i}\int_{s-i\infty}^{s+i\infty}\frac{\Gamma(r+1)}{z^{r+1}}e^{zx}\mathrm{d}z,\quad x>0,$
-provided the integral is well defined. Using $x>0$, we can rewrite the integral as
-$|x|^{r}=x^{r}=\frac{\Gamma(r+1)}{2\pi i}\int_{s-i\infty}^{s+i\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}z=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}t,$ (A.2)
-which is well-defined for all $\mathrm{Re}[r]>-1$ (because $x>0$).
-For $x=0$ the expression is well-defined for all $\mathrm{Re}[r]>0$.
-This completes the proof of the third expression in Lemma 1.
-We proceed with
-$|x|^{r}$ $=x^{r}=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}t+\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-zx}}{z^{r+1}}\mathrm{d}t$
-$=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}+e^{-zx}}{z^{r+1}}\mathrm{d}t,\qquad\qquad\quad x\geq 0,$ (A.3)
-where the second term is zero by Lemma A.2.
+$\int_{C_{R}}f(\zeta)\mathrm{d}\zeta=\int_{C_{R}}e^{i\zeta x}g(\zeta)\mathrm{d}\zeta=\int_{C_{R}}e^{ia\zeta}g(\zeta)\mathrm{d}\zeta,$
+where $a=x>0$ and $g(\zeta)=\frac{1}{(s-i\zeta)^{\nu}}$ is continuous and satisfies $\lim_{|\zeta|\rightarrow\infty}g(\zeta)=0$ for all $\operatorname{Re}(\nu)>0$.
+The second reason for the change of variable is that it enables us to apply Jordan’s lemma with $a>0$, which gives us that $\lim_{R\rightarrow\infty}\int_{C_{R}}e^{ix\zeta}g(\zeta)\mathrm{d}\zeta=0$.
+Thus, by Jordan’s lemma and the Cauchy Integral Theorem, the result now follows.
+For $x=0$ the result follows directly by
+$\int_{-\infty}^{+\infty}\frac{1}{(s+it)^{\nu}}\mathrm{d}t=\left.\frac{i}{\nu-1}\frac{1}{(s+it)^{\nu-1}}\right|_{-\infty}^{+\infty},$
+because the last term is zero for $\operatorname{Re}(\nu)>1$. This completes the proof. $\square$
+Proof of Lemma 2. Consider first the case $x\geq 0$. From the Laplace transformation of $x^{r}$,
+$\int_{0}^{\infty}x^{r}e^{-zx}\mathrm{d}x$ $=\frac{\Gamma(r+1)}{z^{r+1}},\qquad\operatorname{Re}(r)>-1,\quad\operatorname{Re}(z)>0,$ (A.1)
+see e.g.
+Gradshteyn and Ryzhik, (2007, 3.381.4), and the inverse Laplace transform is $x^{r}=\frac{1}{2\pi i}\int_{s-i\infty}^{s+i\infty}\frac{\Gamma(r+1)}{z^{r+1}}e^{zx}\mathrm{d}z$, provided the integral is well defined.
+For $x>0$, we can rewrite the integral as
+$|x|^{r}=x^{r}=\tfrac{\Gamma(r+1)}{2\pi i}\int_{s-i\infty}^{s+i\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}z=\tfrac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}t,$ (A.2)
+which is well-defined for all $\operatorname{Re}(r)>-1$.
+For $x=0$ the expression is well-defined for $\operatorname{Re}(r)>0$.
+This completes the proof of the second expression in Lemma 2.
+Because $\tfrac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-zx}}{z^{r+1}}\mathrm{d}t=0$ for $x>0$ by Lemma 1, we have
+$|x|^{r}=x^{r}$ $=\tfrac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}t+\tfrac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-zx}}{z^{r+1}}\mathrm{d}t$
+$=\tfrac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}+e^{-zx}}{z^{r+1}}\mathrm{d}t,\qquad\qquad\quad x>0.$ (A.3)
 For the case $x<0$, we have the following identity directly from (A.2),
-$|x|^{r}=\left(-x\right)^{r}$ $=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-zx}}{z^{r+1}}\mathrm{d}t,\qquad x<0.$
-Adding $\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}t$ to the right-hand side, which equals zero Lemma A.2, we arrive at
-$|x|^{r}=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-zx}+e^{zx}}{z^{r+1}}\mathrm{d}t,\qquad\text{for }x<0.$ (A.4)
-Combining (A.3) and (A.4) proves the first expression in Lemma 1.
-From (A.2), we also have
-$x^{k}$ $=\qquad\quad\frac{\Gamma(k+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{+zx}}{z^{k+1}}\mathrm{d}t,\quad x>0,$
-$x^{k}$ $=\left(-1\right)^{k}\frac{\Gamma(k+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-zx}}{z^{k+1}}\mathrm{d}t,\quad x<0.$
-This, and $\Gamma(k+1)=k!$ for $k\in\mathbb{N}$, proves the second expression in Lemma 1. This completes the proof. $\square$
-Proof of Theorem 1. The proof is the same as in the proof the Theorem 2 by using the first formula in Lemma 1. $\square$
+$|x|^{r}=\left(-x\right)^{r}$ $=\tfrac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-zx}}{z^{r+1}}\mathrm{d}t,\qquad x<0.$
+Adding $\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}t$ to the right-hand side, which equals zero Lemma 1, we arrive at
+$|x|^{r}=\tfrac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-zx}+e^{zx}}{z^{r+1}}\mathrm{d}t,\qquad\text{for }x<0.$ (A.4)
+Combining (A.3) and (A.4) proves the first expression in Lemma 2 for $x\neq 0$ and $\mathrm{\operatorname{Re}}(r)>-1$.
+For $x=0$ and $\mathrm{\operatorname{Re}}(r)>0$, the left-hand side is $0^{r}=0$ and the integral is zero by Lemma 1.
+For $x=r=0$, the left-hand side is $0^{0}=1$ and the integral simplifies as follows,
+$\tfrac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-zx}+e^{zx}}{z^{r+1}}\mathrm{d}t=\tfrac{1}{2\pi}\int_{-\infty}^{+\infty}\frac{2}{z}\mathrm{d}t=1,$
+where we used (2). This completes the proof of first formula.
+Finally, the last expression in Lemma 2 follows from
+$x^{k}$ $=\qquad\quad\tfrac{\Gamma(k+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{+zx}}{z^{k+1}}\mathrm{d}t,\quad x>0,$
+$x^{k}$ $=\left(-1\right)^{k}\tfrac{\Gamma(k+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-zx}}{z^{k+1}}\mathrm{d}t,\quad x<0,$
+and $\Gamma(k+1)=k!$, where we used (A.2) with $k\in\mathbb{N}$ and $x>0$ and x>0-x>0, respectively. This completes the proof. $\square$
+Proof of Theorem 1. Set $x=X-\xi$ in the first expression of Lemma 2 and take expected value to both sides, then we get
+$\mathbb{E}|X-\xi|^{r}=\tfrac{\Gamma(r+1)}{2\pi}\mathbb{E}\left[\int_{-\infty}^{+\infty}\frac{e^{z\left(X-\xi\right)}+e^{-z\left(X-\xi\right)}}{z^{r+1}}\mathrm{d}t\right],$
+where $z=s+it$ with $s>0$. We consider the cases $\operatorname{Re}(r)>0$ and $\operatorname{Re}(r)\in\left(-1,0\right)$ separately.
+Case $\operatorname{Re}(r)>0$: For all $x\in\mathbb{R}$ we have
+$\mathbb{E}\left[\int_{-\infty}^{+\infty}\left|\frac{e^{z\left(X-\xi\right)}+e^{-z\left(X-\xi\right)}}{z^{r+1}}\right|\mathrm{d}t\right]$
+$=$ $\int_{-\infty}^{+\infty}\mathbb{E}\left|\frac{e^{z\left(X-\xi\right)}+e^{-z\left(X-\xi\right)}}{z^{r+1}}\right|\mathrm{d}t,\quad\text{by Tonelli's theorem}$
+$\leq$ $e^{-s\xi}M_{X}\left(s\right)\int_{-\infty}^{+\infty}\left|\frac{1}{z^{r+1}}\right|\mathrm{d}t+e^{s\xi}M_{X}\left(-s\right)\int_{-\infty}^{+\infty}\left|\frac{1}{z^{r+1}}\right|\mathrm{d}t$
+$\leq$ $e^{\pi|\operatorname{Im}(r)|}\left[e^{-s\xi}M_{X}\left(s\right)+e^{s\xi}M_{X}\left(-s\right)\right]\int_{-\infty}^{+\infty}\frac{1}{\left(s^{2}+t^{2}\right)^{(\operatorname{Re}(r)+1)/2}}\mathrm{d}t$
+$=$ $e^{\pi|\operatorname{Im}(r)|}\left[e^{-s\xi}M_{X}\left(s\right)+e^{s\xi}M_{X}\left(-s\right)\right]\frac{\sqrt{\pi}\Gamma\left(\frac{\operatorname{Re}(r)}{2}\right)}{s^{\operatorname{Re}(r)}\Gamma\left(\frac{\operatorname{Re}(r)+1}{2}\right)}<\infty,$
+where we used $\left|\frac{1}{z^{r+1}}\right|\leq e^{\pi|\operatorname{Im}(r)|}\frac{1}{|z|^{\operatorname{Re}(r)+1}}$.
+Therefore, if $\mathbb{E}e^{\pm sX}<\infty$ and $\operatorname{Re}(r)>0$, then we can interchange the integral and expectations.
+This proves (4) if $\operatorname{Re}(r)>0$.
+Note that ${\rm Pr}\left(X=\xi\right)>0$ is permitted for $\operatorname{Re}(r)>0$.
+Case $\operatorname{Re}(r)\in(-1,0)$.
+We establish the result for $x\neq 0$.
+For this case, we seek an integrable dominating function, $g(x|s,r)$, that satisfies $\sup_{T>0}\left|\int_{-T}^{+T}\frac{e^{zx}+e^{-zx}}{z^{r+1}}\mathrm{d}t\right|\leq g\left(x|s,r\right)$, such that we can interchange the integral and expectation by the dominated convergence theorem.
+First, consider $\int_{-T}^{+T}e^{zx}/z^{r+1}\mathrm{d}t$ (the result for $\int_{-T}^{+T}e^{-zx}/z^{r+1}\mathrm{d}t$ follows similarly). We have
+$\int_{-T}^{+T}\frac{e^{\left(s+it\right)x}}{(s+it)^{r+1}}\mathrm{d}t$ $=e^{sx}I_{T}\left(x|r\right),\quad I_{T}(x|r)\equiv\int_{-T}^{+T}\frac{e^{itx}}{(s+it)^{r+1}}\mathrm{d}t.$
+Let $u=t|x|$, $S=T|x|>0$, and $a=s|x|>0$ . Then
+$I_{T}(x|r)$ $=\int_{-T|x|}^{+T|x|}\frac{e^{iu\cdot\operatorname{sgn}(x)}}{(s+iu/|x|)^{r+1}}\cdot\frac{du}{|x|}=|x|^{r}\int_{-S}^{+S}\frac{e^{iu\operatorname{sgn}(x)}}{(a+iu)^{r+1}}\mathrm{d}u.$
+For $x>0$ (the proof for $x<0$ is similar and $x=0$ is ruled out by assumption), we have
+$\int_{-S}^{+S}\frac{e^{iu}}{(a+iu)^{r+1}}\mathrm{d}u$ $=\int_{0}^{S}\frac{e^{-iu}}{(a-iu)^{r+1}}\mathrm{d}u+\int_{0}^{S}\frac{e^{iu}}{(a+iu)^{r+1}}\mathrm{d}u.$
+It is therefore sufficient to show that there exist a constant, $C_{r}>0$, such that
+$\sup_{S>0,a>0}\left|\int_{0}^{S}\frac{e^{\pm iu}}{(a\pm iu)^{r+1}}\mathrm{d}u\right|\leq\frac{1}{2}C_{r}.$
+Let $U>0$ be an arbitrary positive number. Then for $a>0$, we have
+$\left|\int_{0}^{U}\frac{e^{\pm iu}}{(a\pm iu)^{r+1}}\mathrm{d}u\right|$ $\leq e^{\pi|\operatorname{Im}(r)|}\int_{0}^{U}\frac{1}{|a\pm iu|^{\operatorname{Re}(r)+1}}\mathrm{d}u$
+$\leq e^{\pi|\operatorname{Im}(r)|}\int_{0}^{U}\frac{1}{u^{\operatorname{Re}(r)+1}}\mathrm{d}u$
+$=\frac{e^{\pi|\operatorname{Im}(r)|}}{U^{{\rm Re}\left(r\right)}\left|\operatorname{Re}(r)\right|},\quad\text{for}\quad\operatorname{Re}(r)<0$
+Here we use $\operatorname{Re}(r)<0$ to make sure the last integral is convergent.
+If $S>U$, then integration by parts yields
+$\int_{U}^{S}\frac{e^{\pm iu}}{\left(a\pm iu\right)^{r+1}}{\rm d}u=\pm\left.\frac{e^{iu}}{i(a+iu)^{r+1}}\right|_{\pm U}^{\pm S}+\left(r+1\right)\int_{U}^{S}\frac{e^{\pm iu}}{(a\pm iu)^{r+2}}{\rm d}u,$
+where the last integral has the following bound,
+$\int_{U}^{S}\left|\frac{1}{\left(a\pm iu\right)^{r+2}}\right|du$ $\leq e^{\pi|\operatorname{Im}(r)|}\int_{U}^{S}\frac{1}{u^{\operatorname{Re}(r)+2}}du$
+$=\frac{e^{\pi|\operatorname{Im}(r)|}}{\operatorname{Re}(r)+1}\left(\frac{1}{U^{\operatorname{Re}(r)+1}}-\frac{1}{S^{\operatorname{Re}(r)+1}}\right)$
+$\leq\frac{e^{\pi|\operatorname{Im}(r)|}}{\operatorname{Re}(r)+1}\frac{1}{U^{\operatorname{Re}(r)+1}},\quad\text{for}\quad\operatorname{Re}(r)>-1.$
+So,
+$\left|\int_{U}^{S}\frac{e^{\pm iu}}{\left(a\pm iu\right)^{r+1}}{\rm d}u\right|$ $\leq\left|\frac{1}{(a+iS)^{r+1}}\right|+\left|\frac{1}{(a+iU)^{r+1}}\right|+\left(|r|+1\right)\int_{U}^{S}\left|\frac{1}{\left(a\pm iu\right)^{r+2}}\right|{\rm d}u$
+$\leq e^{\pi|\operatorname{Im}(r)|}\left[\frac{1}{S^{\operatorname{Re}(r)+1}}+\frac{1}{U^{\operatorname{Re}(r)+1}}+\frac{\left(|r|+1\right)}{\operatorname{Re}(r)+1}\frac{1}{U^{\operatorname{Re}(r)+1}}\right]$
+$\leq\frac{e^{\pi|\operatorname{Im}(r)|}}{U^{\operatorname{Re}(r)+1}}\left[2+\frac{|r|+1}{\operatorname{Re}(r)+1}\right],\quad{\rm because\ }S>U>0,$
+and combined we have shown that
+$\left|\int_{0}^{S}\frac{e^{\pm iu}}{(a\pm iu)^{r+1}}\mathrm{d}u\right|$ $\leq\left|\int_{0}^{U}\frac{e^{\pm iu}}{\left(a\pm iu\right)^{r+1}}\mathrm{d}u\right|+\left|\int_{U}^{S}\frac{e^{\pm iu}}{\left(a\pm iu\right)^{r+1}}\mathrm{d}u\right|$
+$\leq\frac{e^{\pi|\operatorname{Im}(r)|}}{U^{{\rm Re}\left(r\right)}\left|\operatorname{Re}(r)\right|}+\frac{e^{\pi|\operatorname{Im}(r)|}}{U^{\operatorname{Re}(r)+1}}\left[2+\frac{|r|+1}{\operatorname{Re}(r)+1}\right]\equiv\frac{1}{2}C_{r},\quad\text{if }S>U.$
+If instead $U\geq S>0$. Then the absolute integral is also bounded by $\frac{1}{2}C_{r}$, because
+$\left|\int_{0}^{S}\frac{e^{\pm iu}}{(a\pm iu)^{r+1}}\mathrm{d}u\right|$ $\leq\int_{0}^{U}\left|\frac{1}{(a\pm iu)^{r+1}}\right|\mathrm{d}u\leq\frac{e^{\pi|\operatorname{Im}(r)|}}{U^{{\rm Re}\left(r\right)}\left|\operatorname{Re}(r)\right|}<\frac{1}{2}C_{r}.$
+We can now conclude that
+$\left|I_{T}(x|r)\right|$ $=|x|^{\operatorname{Re}(r)}\left|\int_{-S}^{S}\frac{e^{iu\operatorname{sgn}(x)}}{(a+iu)^{r+1}}\mathrm{d}u\right|$
+$\leq|x|^{\operatorname{Re}(r)}\left[\left|\int_{0}^{S}\frac{e^{iu}}{(a+iu)^{r+1}}\mathrm{d}u\right|+\left|\int_{0}^{S}\frac{e^{-iu}}{(a-iu)^{r+1}}\mathrm{d}u\right|\right]\leq C_{r}|x|^{\operatorname{Re}(r)},$
+which implies
+$\sup_{T>0}\left|\int_{-T}^{T}\frac{e^{zx}}{z^{r+1}}\mathrm{d}t\right|\leq C_{r}e^{sx}|x|^{\operatorname{Re}(r)},\quad\sup_{T>0}\left|\int_{-T}^{T}\frac{e^{-zx}}{z^{r+1}}dt\right|\leq C_{r}e^{-sx}|x|^{\operatorname{Re}(r)}.$ (A.5)
+Finally, we have
+$\sup_{T>0}\left|\int_{-T}^{+T}\frac{e^{zx}+e^{-zx}}{z^{r+1}}\mathrm{d}t\right|\leq C_{r}\left(e^{sx}+e^{-sx}\right)|x|^{\operatorname{Re}(r)},\quad\operatorname{Re}(r)\in(-1,0),s>0.$
+What remains, is to show that $\mathbb{E}\left(e^{\pm sX}|X|^{p}\right)<\infty$ for $p=\operatorname{Re}(r)\in(-1,0)$. We have
+$\mathbb{E}\left(e^{\pm sX}|X|^{p}\right)$ $=\int_{-\infty}^{+\infty}e^{\pm sX}|X|^{p}{\rm d}F\left(X\right)$
+$=\left[\int_{-\infty}^{-1}e^{\pm sX}|X|^{p}{\rm d}F\left(X\right)+\int_{1}^{+\infty}e^{\pm sX}|X|^{p}{\rm d}F\left(X\right)\right]+\int_{-1}^{1}e^{\pm sX}|X|^{p}{\rm d}F\left(X\right)$
+$\leq\left[\int_{-\infty}^{-1}e^{\pm sX}{\rm d}F\left(X\right)+\int_{1}^{+\infty}e^{\pm sX}{\rm d}F\left(X\right)\right]+e^{s}\int_{-1}^{1}|X|^{p}{\rm d}F\left(X\right)$
+$\leq\mathbb{E}\left(e^{\pm sX}\right)+e^{s}\mathbb{E}|X|^{p}<\infty.$
+Thus both $\mathbb{E}\left(e^{\pm sX}\right)<\infty$ and $\mathbb{E}|X|^{\operatorname{Re}(r)}<\infty$ is needed when $\operatorname{Re}(r)\in(-1,0)$.
+The result now follows by the dominated convergence theorem.
+We have used that $x\neq 0$, such that ${\rm Pr}\left(X=\xi\right)=0$ is needed for $\operatorname{Re}(r)\in(-1,0)$.
+Case $r=0$. In this case $I_{T}(x|s,0)=\int_{-T}^{+T}\frac{e^{itx}}{\left(s+it\right)}\mathrm{d}t$. For $x=0$, we have
+$I_{T}(0|0)=\int_{-T}^{T}\frac{1}{\left(s+it\right)}\mathrm{d}t$ $=\int_{-T}^{T}\frac{s}{s^{2}+t^{2}}\mathrm{d}t-i\int_{-T}^{T}\frac{t}{s^{2}-t^{2}}\mathrm{d}t$
+$=\int_{-T}^{T}\frac{s}{s^{2}+t^{2}}\mathrm{d}t=\arctan\left(\frac{T}{s}\right)-\arctan\left(-\frac{T}{s}\right),$
+which proves that $\left|I_{T}(0|0)\right|\leq\pi.$
+If $x\neq 0$ ($a=s|x|>0$)
+$I_{T}\left(x|0\right)=|x|^{0}\int_{-S}^{+S}\frac{e^{iu\operatorname{sgn}(x)}}{(a+iu)^{0+1}}\mathrm{d}u=\int_{-S}^{+S}\frac{e^{\pm iu}}{a+iu}\mathrm{d}u=2\int_{0}^{S}\frac{a\cos u\pm u\sin u}{a^{2}+u^{2}}\mathrm{d}u,$
+where we used $e^{iu}=\cos u+i\sin u$. For the first term we have
+$\left|\int_{0}^{S}\frac{a\cos u}{a^{2}+u^{2}}du\right|$ $\leq\int_{0}^{S}\frac{a}{a^{2}+u^{2}}du=\arctan\left(\frac{S}{a}\right)\leq\frac{\pi}{2},$
+and for the second term we use $\frac{u}{a^{2}+u^{2}}=\frac{1}{u}-\frac{a^{2}}{u\left(a^{2}+u^{2}\right)}$, to get
+$\left|\int_{0}^{S}\frac{u\sin u}{a^{2}+u^{2}}\mathrm{d}u\right|\leq\left|\int_{0}^{S}\frac{\sin u}{u}\mathrm{d}u\right|+\left|\int_{0}^{S}\frac{a^{2}\sin u}{u\left(a^{2}+u^{2}\right)}\mathrm{d}u\right|.$ (A.6)
+The first term of (A.6) is the Sine integral function, $0\leq\operatorname{Si}(S)\leq\operatorname{Si}(\pi)<\infty$. For the second term in (A.6), we have
+$\left|\int_{0}^{S}\frac{a^{2}\sin u}{u\left(a^{2}+u^{2}\right)}\mathrm{d}u\right|\leq\left|\int_{0}^{\infty}\frac{a^{2}\sin u}{u\left(a^{2}+u^{2}\right)}\mathrm{d}u\right|+\left|\int_{S}^{\infty}\frac{a^{2}\sin u}{u\left(a^{2}+u^{2}\right)}\mathrm{d}u\right|.$ (A.7)
+Next, define
+$J(S;a)=\int_{S}^{\infty}\frac{a^{2}\sin u}{u\left(a^{2}+u^{2}\right)}\mathrm{d}u.$
+Then, for the first term in (A.7), we have the known formula
+$\left|J\left(0;a\right)\right|=J\left(0;a\right)=\int_{0}^{\infty}\frac{a^{2}\sin u}{u\left(a^{2}+u^{2}\right)}\mathrm{d}u=\frac{\pi}{2}e^{-a}\leq\frac{\pi}{2}.$
+For the second term in (A.7) we have $\frac{a^{2}}{\left(a^{2}+u^{2}\right)u}\sin u\geq 0$, for $u\in\left(0,\pi\right]$, whereas the integrand oscillates between positive and negative values when $u>\pi$.
+Define
+$A_{k}^{+}$ $=\int_{2k\pi}^{(2k+1)\pi}\frac{a^{2}\sin u}{\left(a^{2}+u^{2}\right)u}\mathrm{d}u>0,$
+$A_{k}^{-}$ $=\int_{(2k+1)\pi}^{(2k+2)\pi}\frac{a^{2}\sin u}{\left(a^{2}+u^{2}\right)u}\mathrm{d}u<0,$
+then $A_{k}^{+}>|A_{k}^{-}|$, such that $A_{k}^{+}+A_{k}^{-}>0$, because $\frac{a^{2}}{\left(a^{2}+u^{2}\right)u}$ is a decreasing in $u>0$ (the amplitude of oscillation decreases in $u$).
+Now set $K\equiv\lfloor S/2\pi\rfloor$ such that $S\in[2K\pi,2(K+1)\pi)$.
+Then
+$\int_{S}^{2(K+1)\pi}\frac{a^{2}\sin u}{\left(a^{2}+u^{2}\right)u}\mathrm{d}u\leq\int_{2K\pi}^{2(K+1)\pi}\frac{a^{2}\sin u}{\left(a^{2}+u^{2}\right)u}\mathrm{d}u=A_{K}^{+}+A_{K}^{-},$
+and it follows that
+$\left|J(S;a)\right|=J(S;a)\leq\sum_{k=K}^{\infty}(A_{k}^{+}+A_{k}^{-})\leq\sum_{k=0}^{\infty}(A_{k}^{+}+A_{k}^{-})=J\left(0;a\right)\leq\frac{\pi}{2},$
+such that $\sup_{S>0,a>0}\left|\int_{-S}^{+S}\frac{e^{iu}}{a+iu}\mathrm{d}u\right|\leq\frac{\pi}{2}+\operatorname{Si}(\pi)+\frac{\pi}{2}=\pi+\operatorname{Si}(\pi)<\infty$.
+Case ${\rm Re}\left(r\right)=0$, ${\rm Im}\left(r\right)\neq 0$ and $x\neq 0$.
+Proving the identity for the case $r=i\eta$ with $\eta\neq 0$ and $x\neq 0$ turned out to be surprising challenging.
+[^13] While we can handle $x=0$ in the case $r=0$, it is not possible for $r=i\eta$ with $\eta\neq 0$, because $0^{i\eta}$ is not well-defined and the integral $\int_{-T}^{+T}\frac{1}{\left(s+it\right)^{1+i\eta}}$ diverges as $T\rightarrow\infty$.
+Set $S=T|x|$, $u=t|x|$, and $a=s|x|$, For $x>0$ (the proof under $x<0$ is similar), we have
+$\int_{-T}^{+T}\frac{e^{zx}}{z^{r+1}}\mathrm{d}t=|x|^{r}\int_{-S}^{+S}\frac{e^{iu\operatorname{sgn}(x)}}{(a+iu)^{r+1}}\mathrm{d}u=x^{i\eta}\int_{-S}^{+S}f\left(u\right)\mathrm{d}u,\quad f\left(u\right)=\frac{e^{iu}}{(a+iu)^{1+i\eta}}.$
+Since $|x^{i\eta}|=1$, we will show that
+$\sup_{S>0,a>0}\left|\int_{-S}^{+S}f\left(u\right)\mathrm{d}u\right|\leq C_{\eta}.$ (A.8)
+Recall the upper incomplete Gamma function,
+$\Gamma\left(\sigma,z\right)\equiv\int_{z}^{\infty+i{\rm Im}\left(z\right)}t^{\sigma-1}e^{-t}\mathrm{d}t=e^{-z}\int_{0}^{\infty}\frac{e^{-u}}{(z+u)^{1-\sigma}}\mathrm{d}u,\quad{\rm with}\ t=z+u$
+and with $z=\mp iS-a$, $\sigma=-i\eta$, we have
+$e^{-z}\int_{0}^{\infty}\frac{e^{-u}}{(z+u)^{1-\sigma}}\mathrm{d}u$ $=e^{a}\int_{0}^{\infty}\frac{e^{-\left(u\mp iS\right)}}{(u\mp iS-a)^{1+i\eta}}\mathrm{d}u$
+$=-ie^{a}\int_{\pm S}^{\pm S+i\infty}\frac{e^{i\tau}}{(-i\tau-a)^{1+i\eta}}\mathrm{d}\tau,\quad{\rm with}\ \tau=\pm S+iu$
+$=ie^{a-\pi\eta}\int_{\pm S}^{\pm S+i\infty}f\left(\tau\right)\mathrm{d}\tau.$
+and then if we can show
+$\int_{\pm S}^{\pm S+i\infty}f\left(\tau\right)\mathrm{d}\tau=\int_{\pm S}^{\pm\infty}f\left(\tau\right)\mathrm{d}\tau,$ (A.9)
+then we have $\int_{\pm S}^{\pm\infty}f\left(\tau\right)\mathrm{d}\tau=-ie^{\pi\eta-z}\Gamma\left(\sigma,z\right)$. To proof (A.9), we construct the following integral paths,
+$\int_{\pm K}^{\pm S}f\left(\tau\right)\mathrm{d}\tau+\int_{\pm S}^{\pm S+iK}f\left(\tau\right)\mathrm{d}\tau+\int_{\pm S+iK}^{\pm K}f\left(\tau\right)\mathrm{d}\tau=0$
+and $f\left(\tau\right)$ is an analytic function in the constructed closed contour (given $S\neq 0$), so according to Cauchy integral theorem, the integral above is zero.
+So to proof (A.9), we need to show the last term is zero when $K\rightarrow+\infty$.
+Define a semi-arc with length $R$, starting from $\pm S+iK$ with $\theta=\pi/2$ to $\pm K$ with $\theta=0$ for +S+S and $\theta=\pi$ for S-S.
+Take +S+S for example (S-S is the same), we have $u=Re^{i\theta}=R\left(\cos\theta+i\sin\theta\right)$, $\theta=\left[0,\frac{\pi}{2}\right]$, $du=iRe^{i\theta}d\theta$, such that
+$\left|\int_{S+iK}^{K}\frac{e^{iu}}{(a+iu)^{1+i\eta}}du\right|$ $=\left|\int_{\pi/2}^{0}\frac{e^{iR\cos\theta-R\sin\theta}}{(a+Re^{i\theta})^{1+i\eta}}iRe^{i\theta}d\theta\right|$
+$\leq\int_{0}^{\pi/2}\frac{e^{-R\sin\theta}}{\left|(a+Re^{i\theta})^{1+i\eta}\right|}Rd\theta$
+$\leq e^{\left|\eta\right|\pi}\int_{0}^{\pi/2}\frac{Re^{-R\sin\theta}}{\left|a-R\right|}d\theta$
+$\leq e^{\left|\eta\right|\pi}\int_{0}^{\pi/2}\frac{Re^{-R\frac{2}{\pi}\theta}}{\left|a-R\right|}d\theta$
+$=e^{\left|\eta\right|\pi}\frac{R}{\left|a-R\right|}\frac{\pi}{2R}\left(1-e^{-R}\right)$
+$=0\quad{\rm when}\quad R\rightarrow\infty$
+where we use $\sin\theta\geq\frac{2}{\pi}\theta$, for $\theta\in\left[0,\frac{\pi}{2}\right]$ , $\sin\theta\geq\frac{2}{\pi}(\pi-\theta)$ for $\theta\in\left[\frac{\pi}{2},\pi\right]$ and $\left|\frac{1}{z^{1+i\eta}}\right|\leq\frac{e^{\pi|\eta|}}{|z|}\leq\frac{e^{\pi|\eta|}}{|a-R|}$ where $z=a+Re^{i\theta}$.We have
+$\int_{-S}^{+S}\frac{e^{iu}}{\left(a+iu\right)^{1+i\eta}}\mathrm{d}u$ $=\int_{-\infty}^{+\infty}\frac{e^{iu}}{\left(a+iu\right)^{1+i\eta}}\mathrm{d}u-\int_{S}^{+\infty}\frac{e^{iu}}{\left(a+iu\right)^{1+i\eta}}\mathrm{d}u-\int_{-\infty}^{-S}\frac{e^{iu}}{\left(a+iu\right)^{1+i\eta}}\mathrm{d}u$
+$=\frac{2\pi e^{-a}}{\Gamma(1+i\eta)}+ie^{\pi\eta-a}\Gamma(-i\eta,-iS-a)-ie^{-\pi\eta-a}\Gamma(-i\eta,iS-a),$
+where the first integral follows from Lemma 2 (set $z=a+iu$ and $x=1$). So, by defining $F\left(S,a\right)$ function as
+$F\left(S,a\right)=e^{-a}\left|\Gamma(-i\eta,iS-a)\right|$ (A.10)
+we have
+$\left|\int_{-S}^{+S}\frac{e^{iu}}{\left(a+iu\right)^{1+i\eta}}\mathrm{d}u\right|\leq\frac{2\pi e^{-a}}{\left|\Gamma(1+i\eta)\right|}+e^{\pi\eta}F\left(-S,a\right)+e^{-\pi\eta}F\left(S,a\right).$ (A.11)
+With $\sigma=-i\eta$ and $z=\pm iS-a$, we have
+$F\left(\pm S,a\right)=e^{-a}\left|\Gamma(-i\eta,\pm iS-a)\right|=\left|\int_{0}^{\infty}\frac{e^{-u}}{(z+u)^{1-\sigma}}\mathrm{d}u\right|.$
+Now, if $|S|>1$, then $\left|z+u\right|=\sqrt{\left(u-a\right)^{2}+S^{2}}>1$, such that $\left|(z+u)^{-i\eta-1}\right|\leq\left|z+u\right|^{-1}e^{|\eta|\pi}\leq e^{|\eta|\pi}$, we have the inequality,
+$F\left(\pm S,a\right)\leq e^{\pi|\eta|}\left|\int_{0}^{\infty}e^{-u}\mathrm{d}u\right|=e^{\pi|\eta|}.\quad|S|>1$ (A.12)
+If instead, $|S|\leq 1$, then we can use, $\Gamma(\sigma,z)=\Gamma(\sigma)-\gamma(\sigma,z)$, where
+$\gamma(\sigma,z)=\int_{0}^{z}t^{\sigma-1}e^{-t}\mathrm{d}t=\frac{z^{\sigma}}{\sigma}+z^{\sigma}\int_{0}^{1}u^{\sigma-1}\left(e^{-zu}-1\right)\mathrm{d}u,$ (A.13)
+is the lower incomplete Gamma function.
+Set $z=\pm iS-a$, and $\sigma=-i\eta$ where $\eta\neq 0$, with $\left|z^{\sigma}\right|=\left|(-a\pm iS)^{-i\eta}\right|\leq e^{|\eta|\pi}$ and $\left|\frac{z^{\sigma}}{\sigma}\right|\leq\frac{e^{|\eta|\pi}}{|\eta|}$, we have
+$\left|\gamma(-i\eta,\pm iS-a)\right|$ $\leq\frac{e^{|\eta|\pi}}{|\eta|}+e^{|\eta|\pi}\int_{0^{+}}^{1}\left|u^{-i\eta-1}\left(e^{-\left(\pm iS-a\right)u}-1\right)\right|\mathrm{d}u$ (A.14)
+And because $\sigma=-i\eta$ where $\eta\neq 0$, $u^{\sigma-1}$ is not well-define when $u=0$.
+Therefore, the integral in the right-hand side is starting from $0^{+}$ to 1, and we only consider the case when $u\in\left(0,1\right]$ below.
+We now focus on $g\left(u\right)$ given by+
+$g\left(u\right)=e^{-(\pm iS-a)u},$
+with $g\left(0\right)=1$, $g^{\prime}\left(u\right)=-g\left(u\right)(\pm iS-a)$ such that $g^{\prime}\left(0\right)=-(\pm iS-a)$, and $g^{\prime\prime}\left(u\right)=g\left(u\right)(\pm iS-a)^{2}$.
+By Taylor’s Theorem, for some $\xi\in\left[0,u\right]$, we have
+$g\left(u\right)=1-(\pm iS-a)u+\tfrac{1}{2}(\pm iS-a)^{2}g\left(\xi\right)u^{2}$
+with $\left|g\left(\xi\right)\right|=e^{a\xi}\leq e^{au}$, we have
+$e^{-a}u^{-1}\left|g\left(u\right)-1\right|\leq e^{-a}\left|\pm iS-a\right|+\tfrac{1}{2}(\pm iS-a)^{2}e^{a\left(u-1\right)}u.\quad u\in\left(0,1\right]$ (A.15)
+where for the first term in the last expression above, we have,
+$e^{-a}|\pm iS-a|=e^{-a}\sqrt{a^{2}+S^{2}}=\sqrt{\left(ae^{-a}\right)^{2}+\left(Se^{-a}\right)^{2}}\leq\tfrac{1}{e}+|S|,$ (A.16)
+where we use $ae^{-a}\leq\frac{1}{e}$ for any $a\geq 0$. As for the second term, we have
+$|\pm iS-a|^{2}e^{a\left(u-1\right)}=e^{a\left(u-1\right)}\left(a^{2}+S^{2}\right)\leq\frac{4}{e^{2}(1-u)^{2}}+S^{2}$ (A.17)
+Define the function $f_{a}\left(u\right)$ as
+$f_{a}\left(u\right)=e^{-a}\left|u^{\sigma-1}\left(g\left(u\right)-1\right)\right|.\quad u\in\left(0,1\right]$ (A.18)
+and from (A.15), (A.16) and (A.17), we have
+$f_{a}\left(u\right)\leq u^{-1}e^{-a}\left|g\left(u\right)-1\right|\leq\tfrac{1}{e}+|S|+\tfrac{1}{2}\left(\frac{4}{e^{2}(1-u)^{2}}+S^{2}\right)u.$ (A.19)
+Then for any given $\delta\in\left(0,1\right)$, for all $u\in\left(0,\delta\right)$, we have
+$\left|\int_{0^{+}}^{\delta}f_{a}(u)\mathrm{d}u\right|$ $\leq\int_{0^{+}}^{\delta}\left(\frac{1}{e}+|S|+\frac{1}{2}\left(\frac{4}{e^{2}(1-\delta)^{2}}+S^{2}\right)u\right)\mathrm{d}u$
+$=\delta\left(\frac{1}{e}+|S|\right)+\frac{\delta^{2}}{4}\left(\frac{4}{e^{2}(1-\delta)^{2}}+S^{2}\right)$
+$\leq\delta\left(\frac{1}{e}+1\right)+\frac{\delta^{2}}{4}\left(\frac{4}{e^{2}(1-\delta)^{2}}+1\right),\quad|S|\leq 1.$
+Next, for $u\in\left[\delta,1\right]$, we have
+$f_{a}\left(u\right)=u^{-1}e^{-a}\left|e^{-(\pm iS-a)u}-1\right|\leq\tfrac{1}{\delta}e^{-a}\left(e^{au}+1\right)\leq\tfrac{1}{\delta}\left(e^{-a}+1\right)\leq\tfrac{2}{\delta}$
+such that for $|S|\leq 1$, we have
+$\left|\int_{0^{+}}^{1}f_{a}(u)\mathrm{d}u\right|$ $\leq\left|\int_{0^{+}}^{\delta}f_{a}(u)\mathrm{d}u\right|+\left|\int_{\delta}^{1}f_{a}(u)\mathrm{d}u\right|$
+$\leq\delta\left(\frac{1}{e}+1\right)+\frac{\delta^{2}}{4}\left(\frac{4}{e^{2}(1-\delta)^{2}}+1\right)+\tfrac{2}{\delta}\left(1-\delta\right)$
+$=\tfrac{1}{2}\left(\tfrac{1}{e}+1\right)+\left(\tfrac{1}{e^{2}}+\tfrac{1}{16}\right)+2,\quad\delta=\tfrac{1}{2}$
+Recall from (A.14) and the definition of $f_{a}(u)$ in (A.18), we have for $|S|\leq 1$
+$e^{-a}\left|\gamma(-i\eta,\pm iS-a)\right|$ $\leq e^{-a}\frac{e^{|\eta|\pi}}{|\eta|}+e^{|\eta|\pi}\int_{0^{+}}^{1}f_{a}(u)\mathrm{d}u$
+$\leq e^{|\eta|\pi}\tfrac{1}{|\eta|}+e^{|\eta|\pi}\left[\tfrac{1}{2}\left(\tfrac{1}{e}+1\right)+\left(\tfrac{1}{e^{2}}+\tfrac{1}{16}\right)+2\right]=C_{\eta}^{(1)}$
+such that for $|S|\leq 1$, we have
+$F\left(\pm S,a\right)$ $=e^{-a}\left|\Gamma(-i\eta,\pm iS-a)\right|$
+$\leq e^{-a}\left[\left|\Gamma(-i\eta)\right|+\left|\gamma(-i\eta,\pm iS-a)\right|\right]$
+$\leq\left|\Gamma(-i\eta)\right|+C_{\eta}^{(1)}$
+Combined with the case when $\left|S\right|>1$ in (A.12), we have
+$F\left(\pm S,a\right)\leq\max\left(e^{\pi|\eta|},\left|\Gamma(-i\eta)\right|+C_{\eta}^{(1)}\right)=C_{\eta}^{(2)}$
+Finally, from (A.11), we have
+$\left|\int_{-S}^{+S}\frac{e^{iu}}{\left(a+iu\right)^{1+i\eta}}\mathrm{d}u\right|\leq\frac{2\pi}{\left|\Gamma(1+i\eta)\right|}+\left(e^{\pi\eta}+e^{-\pi\eta}\right)C_{\eta}^{(2)}=C_{\eta}.$
+$\square$
 Proof of Theorem 2.
-We prove the result for discrete and continuous distributions separately.
-Consider first a non-negative discrete random variable, $X$, with support $\mathcal{X}$.
-Rewriting $\mathbb{E}[X^{r}]=\sum_{x\in\mathcal{X}}x^{r}p(x)$ using the third formula in Lemma 1 yields
-$\mathbb{E}[X^{r}]$ $=\frac{\Gamma(r+1)}{2\pi}\sum_{x\in\mathcal{X}}\left[\int_{-\infty}^{\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}t\right]p(x)$
-$=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{\infty}\sum_{x\in\mathcal{X}}\frac{e^{zx}}{z^{r+1}}p(x)\mathrm{d}t$
-$=\frac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{\infty}\frac{M_{X}\left(z\right)}{z^{r+1}}\mathrm{d}t,$
-for $r+1\in\mathbb{C}^{+}$ if $\Pr(X=0)=0$ and $r\in\mathbb{C}^{+}$ if $\Pr(X=0)>0$. For a continuous distribution, the same line of arguments gives us,
-$\mathbb{E}[X^{r}]=\tfrac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{\infty}\left[\int_{-\infty}^{\infty}\tfrac{e^{zx}}{z^{r+1}}\mathrm{d}t\right]f(x)\mathrm{d}x=\tfrac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{\infty}\tfrac{M_{X}\left(u\right)}{z^{r+1}}\mathrm{d}t.$
-This completes the proof. $\square$
-Proof of Theorem 3. The proof is the same as in the proof the Theorem 1 by using the second formula in Lemma 1. $\square$
+From Lemma 2 we have $x^{r}=\tfrac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}t,$ for $x>0$ and ${\rm Re}\left(r\right)>-1$.
+With the convention $x_{+}^{r}\equiv x^{r}\mathds{1}_{\{x>0\}}$ we have
+$x_{+}^{r}=\tfrac{\Gamma(r+1)}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{zx}}{z^{r+1}}\mathrm{d}t,\quad x\neq 0,\operatorname{Re}(r)>-1,$
+because $\int_{-\infty}^{+\infty}\frac{e^{itx}}{z^{r+1}}=0$ for $x<0$ by Lemma 1, which also gives us that the identity holds for $x=0$ and $\operatorname{Re}(r)>0$.
+When $\operatorname{Re}(r)>0$, we also use Fubini’s theorem, such that
+$\mathbb{E}\left[\int_{-\infty}^{+\infty}\left|\frac{e^{z\left(X-\xi\right)}}{z^{r+1}}\right|\mathrm{d}t\right]\leq e^{\pi|{\rm Im}\left(r\right)|}e^{-s\xi}\mathbb{E}\left[e^{sX}\right]\frac{\sqrt{\pi}\Gamma\left(\frac{\operatorname{Re}(r)}{2}\right)}{s^{\operatorname{Re}(r)}\Gamma\left(\frac{\operatorname{Re}(r)+1}{2}\right)}<\infty$
+Therefore, if $\mathbb{E}e^{sX}<\infty$ and $\operatorname{Re}(r)>0$, then we can interchange the integral and expectations.
+This proves (6) if $\operatorname{Re}(r)>0$.
+Note that ${\rm Pr}\left(X=\xi\right)>0$ is permitted for $\operatorname{Re}(r)>0$.
+When $\operatorname{Re}(r)\in(-1,0)$, from $\sup_{T>0}\left|\int_{-T}^{T}\frac{e^{zx}}{z^{r+1}}\mathrm{d}t\right|\leq C_{r}e^{sx}|x|^{\operatorname{Re}(r)}$ in (A.5), we need both the existence of $\mathbb{E}e^{sX}<\infty$ and the expectation of absolute moments.
+For $\operatorname{Re}(r)=0$, the proof of Theorem 1 still applied.
+$\square$
+Proof of Corollary 1. Follows from Theorem 2. Can also be shown directly by setting $x=X-\xi$ in the second equation of Lemma 2 and take expected value to both sides.
+$\square$
+Proof of Theorem 3. Set $x=X-\xi$ in the third expression of Lemma 2 and take expected value to both sides. $\square$
 Appendix B Result for the Dynamic Models
-B.1 MGF for ARG model
+The MGF in Heston-Nandi GARCH Model
+Proof of Proposition 1. We first have
+$\mathbb{E}_{t}\left[\exp\left(v_{1}r_{t+1}+v_{2}h_{t+2}\right)\right]=$ $\exp\left[v_{1}r+v_{2}\omega+\left[v_{1}\left(\lambda-\tfrac{1}{2}\right)+v_{2}\left(\beta+\alpha\gamma^{2}\right)\right]h_{t+1}\right]$
+$\times\mathbb{E}_{t}\left[v_{2}\alpha z_{t+1}^{2}+\left(v_{1}-2v_{2}\alpha\gamma\right)\sqrt{h_{t+1}}z_{t+1}\right]$
+$=$ $\exp\left[v_{1}r+v_{2}\omega+\left[v_{1}\left(\lambda-\tfrac{1}{2}\right)+v_{2}\left(\beta+\alpha\gamma^{2}\right)\right]h_{t+1}\right$.
+$\left.-\frac{1}{2}\log\left(1-2v_{2}\alpha\right)+\frac{\left(v_{1}-2v_{2}\alpha\gamma\right)^{2}}{2\left(1-2v_{2}\alpha\right)}h_{t+1}\right]$
+$=$ $\exp\left[v_{1}r+v_{2}\omega-\tfrac{1}{2}\log\left(1-2v_{2}\alpha\right)\right$.
+$\left.+\left(v_{1}\left(\lambda-\tfrac{1}{2}\right)+v_{2}\left(\beta+\alpha\gamma^{2}\right)+\frac{\left(v_{1}-2v_{2}\alpha\gamma\right)^{2}}{2\left(1-2v_{2}\alpha\right)}\right)h_{t+1}\right].$
+Next, we show by induction that the conditional MGF has the form,
+$\mathbb{E}_{t}\left(\exp\left(s\sum_{i=1}^{H}r_{t+i}\right)\right)=\exp\left(A(H,s)+B(H,s)h_{t+1}\right).$
+For $h=1$ we have
+$\mathbb{E}_{t}\left(\exp\left(sr_{t+1}\right)\right)$ $=\exp\left(sr+\left(s\left(\lambda-\tfrac{1}{2}\right)+\tfrac{s^{2}}{2}\right)h_{t+1}\right)=\exp\left(A(1,s)+B(1,s)h_{t+1}\right)$
+such that $A(1,s)=sr$ and $B(1,s)=s\left(\lambda-\tfrac{1}{2}\right)+\frac{s^{2}}{2}$. Next, suppose that the MGF has an affine structure for $h$, then for $h+1$ we find
+$\mathbb{E}_{t}\left(\exp\left(s\sum_{i=1}^{h+1}r_{t+i}\right)\right)$ $=\mathbb{E}_{t}\left[\mathbb{E}_{t+1}\left(\exp\left(s\sum_{i=1}^{h+1}r_{t+i}\right)\right)\right]$
+$=\mathbb{E}_{t}\left[\exp\left(sR_{t+1}\right)\mathbb{E}_{t+1}\left(\exp\left(s\sum_{i=1}^{h}r_{t+1+i}\right)\right)\right]$
+$=\mathbb{E}_{t}\left[\exp\left(sR_{t+1}+A(h,s)+B(h,s)h_{t+2}\right)\right]$
+$=\exp\left(A(h,s)+sr+B(h,s)\omega-\frac{1}{2}\log\left(1-2B(h,s)\alpha\right)\right)$
+$\quad\times\exp\left(\left(s\left(\lambda-\frac{1}{2}\right)+B(h,s)\left(\beta+\alpha\gamma^{2}\right)+\frac{\left(s-2B(h,s)\alpha\gamma\right)^{2}}{2\left(1-2B(h,s)\alpha\right)}\right)h_{t+1}\right)$
+$=\exp\left(A(h+1,s)+B(h+1,s)h_{t+1}\right),$
+where
+$A(h+1,s)$ $=A(h,s)+sr+B(h,s)\omega-\frac{1}{2}\log\left(1-2B(h,s)\alpha\right)$
+$B(h+1,s)$ $=s\left(\lambda-\frac{1}{2}\right)+B(h,s)\left(\beta+\alpha\gamma^{2}\right)+\frac{\left(s-2B(h,s)\alpha\gamma\right)^{2}}{2\left(1-2B(h,s)\alpha\right)},$
+for any sufficiently small $s=\operatorname{Re}(z)>0$, such that $B(h,s)<\frac{1}{2\alpha}$ for all $h=1,\ldots,H$, to ensure that $\log\left(1-2B(h,s)\alpha\right)\in\mathbb{R}$.
+This completes the proof.
+$\square$
+MGF for ARG model
 Proof of Proposition 2. First, the dynamic process of $\theta_{t}$ can be rewritten as
@@
 $=\exp\left(A(h,z)+\sum_{j=1}^{p-1}B_{j+1}(h,z)X_{T+1-j}\right)$
-×$\quad\times\exp\left(-\delta\log\left(1-\eta B_{1}(h,z)\right)+\frac{\eta B_{1}(h,z)}{1-\eta B_{1}(h,z)}\sum_{j=1}^{p}\beta_{j}X_{T+1-j}\right)$
+$\quad\times\exp\left(-\delta\log\left(1-\eta B_{1}(h,z)\right)+\frac{\eta B_{1}(h,z)}{1-\eta B_{1}(h,z)}\sum_{j=1}^{p}\beta_{j}X_{T+1-j}\right)$
 $=\exp\left(A(h+1,z)+\sum_{j=1}^{p}B_{j}(h+1,z)x_{T+1-j}\right),$
@@
 $\square$
-B.2 The MGF in Heston-Nandi GARCH Model
-We first have
-$\mathbb{E}_{t}\left[\exp\left(v_{1}r_{t+1}+v_{2}h_{t+2}\right)\right]=$ $\exp\left[v_{1}r+v_{2}\omega+\left[v_{1}\left(\lambda-\tfrac{1}{2}\right)+v_{2}\left(\beta+\alpha\gamma^{2}\right)\right]h_{t+1}\right]$
-×$\times\mathbb{E}_{t}\left[v_{2}\alpha z_{t+1}^{2}+\left(v_{1}-2v_{2}\alpha\gamma\right)\sqrt{h_{t+1}}z_{t+1}\right]$
-$=$ $\exp\left[v_{1}r+v_{2}\omega+\left[v_{1}\left(\lambda-\tfrac{1}{2}\right)+v_{2}\left(\beta+\alpha\gamma^{2}\right)\right]h_{t+1}\right.$
-$\left.-\frac{1}{2}\log\left(1-2v_{2}\alpha\right)+\frac{\left(v_{1}-2v_{2}\alpha\gamma\right)^{2}}{2\left(1-2v_{2}\alpha\right)}h_{t+1}\right]$
-$=$ $\exp\left[v_{1}r+v_{2}\omega-\tfrac{1}{2}\log\left(1-2v_{2}\alpha\right)\right.$
-$\left.+\left(v_{1}\left(\lambda-\tfrac{1}{2}\right)+v_{2}\left(\beta+\alpha\gamma^{2}\right)+\frac{\left(v_{1}-2v_{2}\alpha\gamma\right)^{2}}{2\left(1-2v_{2}\alpha\right)}\right)h_{t+1}\right].$
-Next, we show by induction that the conditional MGF has the form,
-$\mathbb{E}_{t}\left(\exp\left(s\sum_{i=1}^{H}r_{t+i}\right)\right)=\exp\left(A(H,s)+B(H,s)h_{t+1}\right).$
-For $h=1$ we have
-$\mathbb{E}_{t}\left(\exp\left(sr_{t+1}\right)\right)$ $=\exp\left(sr+\left(s\left(\lambda-\tfrac{1}{2}\right)+\tfrac{s^{2}}{2}\right)h_{t+1}\right)=\exp\left(A(1,s)+B(1,s)h_{t+1}\right)$
-such that $A(1,s)=sr$ and $B(1,s)=s\left(\lambda-\tfrac{1}{2}\right)+\frac{s^{2}}{2}$. Next, suppose that the MGF has an affine structure for $h$, then for $h+1$ we find
-$\mathbb{E}_{t}\left(\exp\left(s\sum_{i=1}^{h+1}r_{t+i}\right)\right)$ $=\mathbb{E}_{t}\left[\mathbb{E}_{t+1}\left(\exp\left(s\sum_{i=1}^{h+1}r_{t+i}\right)\right)\right]$
-$=\mathbb{E}_{t}\left[\exp\left(sR_{t+1}\right)\mathbb{E}_{t+1}\left(\exp\left(s\sum_{i=1}^{h}r_{t+1+i}\right)\right)\right]$
-$=\mathbb{E}_{t}\left[\exp\left(sR_{t+1}+A(h,s)+B(h,s)h_{t+2}\right)\right]$
-$=\exp\left(A(h,s)+sr+B(h,s)\omega-\frac{1}{2}\log\left(1-2B(h,s)\alpha\right)\right)$
-×$\quad\times\exp\left(\left(s\left(\lambda-\frac{1}{2}\right)+B(h,s)\left(\beta+\alpha\gamma^{2}\right)+\frac{\left(s-2B(h,s)\alpha\gamma\right)^{2}}{2\left(1-2B(h,s)\alpha\right)}\right)h_{t+1}\right)$
-$=\exp\left(A(h+1,s)+B(h+1,s)h_{t+1}\right),$
-where
-$A(h+1,s)$ $=A(h,s)+sr+B(h,s)\omega-\frac{1}{2}\log\left(1-2B(h,s)\alpha\right)$
-$B(h+1,s)$ $=s\left(\lambda-\frac{1}{2}\right)+B(h,s)\left(\beta+\alpha\gamma^{2}\right)+\frac{\left(s-2B(h,s)\alpha\gamma\right)^{2}}{2\left(1-2B(h,s)\alpha\right)},$
-for any sufficiently small $s=\operatorname{Re}(z)>0$, such that $B(h,s)<\frac{1}{2\alpha}$ for all $h=1,\ldots,H$, to ensure that $\log\left(1-2B(h,s)\alpha\right)\in\mathbb{R}$.
-This completes the proof.
-$\square$
 Supplement
-An Direct Proof of Lemma A.1.
-Proof of Lemma A.1. By the substitution, $y=zx\in\mathbb{C}$, we have
-$\lim_{K\rightarrow\infty}\int_{0}^{K}x^{r}e^{-zx}\mathrm{d}x=\frac{1}{z^{r+1}}\int_{0}^{sK+itK}y^{r}e^{-y}\mathrm{d}y.$
-The function $f(y)\equiv y^{r}e^{-y}$, has a pole at $y=0$ if $r<0$, and is a multi-valued function when $r$ is a non-integer.
-We will adopt the principal value by restricting the angle of $y=\lambda e^{i\theta}$ to $\theta\in(-\pi,\pi]$, and construct a closed contour integral that avoids the pole at $y=0$.
-The contour is illustrated in Figure S.1.
-Figure S.1: The closed contour for the integral, $\oint y^{r}e^{-y}\mathrm{d}y$.
-The function, $f(y)$, is analytic over the closed contour for $\lambda>0$, (i.e. it does not include any singularity). Therefore, by the Cauchy Integral Theorem, we have
-$0=\oint f(y)\mathrm{d}y$ $=\int_{AB}f(y)\mathrm{d}y+\int_{BC}f(y)\mathrm{d}y+\int_{CD}f(y)\mathrm{d}y+\int_{DA}f(y)\mathrm{d}y,$
-where the integral we seek is $-\int_{CD}f(y)\mathrm{d}y$ as $\lambda\rightarrow 0$ and $K\rightarrow\infty$.
-For the integral over the arc, DA, we have $y=\lambda e^{i\theta}=\lambda\left(\cos\theta+i\sin\theta\right)$, and $dy=i\lambda e^{i\theta}d\theta$, such that
-$\left|I_{DA}\right|=\left|\int_{DA}y^{r}e^{-y}\mathrm{d}y\right|$ $=\left|\int_{\pi/2}^{0}\lambda^{r}e^{i\theta r}e^{-\lambda\left(\cos\theta+i\sin\theta\right)}i\lambda e^{i\theta}\mathrm{d}\theta\right|$
-$\leq\left|\lambda^{r+1}\right|e^{|\lambda|}\left|\int_{\pi/2}^{0}e^{i\theta r}e^{-i\lambda\sin\theta}ie^{i\theta}\mathrm{d}\theta\right|$
-$\leq\left|\lambda^{r+1}\right|e^{|\lambda|}\left|\int_{\pi/2}^{0}\mathrm{d}\theta\right|=\tfrac{\pi}{2}\lambda^{{\rm Re}\left(r\right)+1}e^{|\lambda|}\rightarrow 0,\quad\text{as }\lambda\rightarrow 0\text{ for }{\rm Re}\left(r\right)>-1.$
-For the BC-integral, we have $y=sK+iuK$ and $\mathrm{d}y=iK\mathrm{d}u$, such that
-$\left|I_{BC}\right|=\left|\int_{sK}^{sK+itK}y^{r}e^{-y}\mathrm{d}y\right|$ $=\left|iK\int_{0}^{t}\left(sK+iuK\right)^{r}e^{-\left(sK+iuK\right)}\mathrm{d}u\right|$
-$=\left|iK^{r+1}e^{-sK}\int_{0}^{t}\left(s+iu\right)^{r}e^{-iuK}\mathrm{d}u\right|$
-$\leq K^{{\rm Re}\left(r\right)+1}e^{-sK}\left|\int_{0}^{t}\left|\left(s+iu\right)^{r}\right|\mathrm{d}u\right|\rightarrow 0,\qquad\text{as }K\rightarrow\infty,$
-because $e^{-sK}$ decays faster than $K^{r+1}$ increases. So, as $K\rightarrow\infty$ and $\lambda\rightarrow 0$, we have
-$0=\lim_{K\rightarrow\infty}\lim_{\lambda\rightarrow 0}\left(\int_{AB}f(y)\mathrm{d}y+\int_{CD}f(y)\mathrm{d}y\right)=\lim_{K\rightarrow\infty}\left(\int_{0}^{sK}f(y)\mathrm{d}y+\int_{sK+itK}^{0}f(y)\mathrm{d}y\right).$
-The integral we seek is therefore given by,
-$\lim_{K\rightarrow\infty}\int_{0}^{sK+itK}f(y)\mathrm{d}y=\lim_{K\rightarrow\infty}\int_{0}^{sK}f(y)\mathrm{d}y=\Gamma(r+1).$
-This completes the proof. $\square$
+Equation A.1.
+A proof of Consider the case where $z$ is real and positive, i.e.
+$z=s>0$.
+From the definition of the gamma function and the substitution, $y\mapsto sx$, we have $\Gamma(r+1)=\int_{0}^{\infty}y^{r}e^{-y}\mathrm{d}y=\int_{0}^{\infty}s^{r}x^{r}e^{-sx}s\mathrm{d}x$, which proves the identity for any $z\in\mathbb{R}^{+}$.
+Both sides of (A.1) are holomorphic functions for $\operatorname{Re}(z)>0$, and since they agree on $z\in\mathbb{R}^{+}$, it follows by the identity theorem that they agree for all $\operatorname{Re}(z)>0$.
 NIG Moments
@@
 For strictly positive variables we have the following expression for negative fractional moments, $r<0$,
-[…]
+$\mathbb{E}[X^{r}]=\frac{1}{\Gamma(-r)}\int_{0}^{\infty}\frac{1}{t^{r+1}}M_{X}(-t)dt,\qquad(X>0),$
 see Schürger, (2002, theorem 1.1).
-Code for Implementation
+Some Computer Code
 Moments in HNG Model